### v2

# On Long-Term Problems in Multiplicative Ideal Theory and Factorization Theory

Abstract.
In this survey article we discuss key open problems which could serve as a guidance for further research directions of multiplicative ideal theory and factorization theory.
Key words and phrases:
multiplicative ideal theory, factorization theory, Prüfer domains, Gorenstein homological algebra, Mori monoids, Krull monoids
2020 Mathematics Subject Classification:
Primary 13A05, 13A15;
Secondary 13F05, 13F15, 13D22, 13D30 20M12, 20M13
The first author was supported by the Austrian Science Fund FWF, Project P36852-N
The second author was supported in part by the Basic Science Research Program through the National Research Foundation of Korea (NRF), funded by the Ministry of Education (2021R1I1A3047469).
1. Introduction
The origins of multiplicative ideal theory can be traced back to the beginnings of commutative algebra in the first half of the 20th century. We refer to articles of Halter-Koch [84, 88] for a discussion of the early development. The monographs of Gilmer [77] and Larsen-McCarthy [107], which appeared in the 1970s, strongly influenced the area. Building on the work of Lorenzen and Aubert, Halter-Koch showed (in his book [84], published 1998) that large parts of multiplicative ideal theory can be derived in the purely multiplicative setting of commutative monoids. This viewpoint also turned out to be crucial for the further development of factorization theory.
Factorization theory has its origin in algebraic number theory. Starting points were results (in the 1960s) of Carlitz and of Narkiewicz, who studied the asymptotic behavior of arithmetic counting functions in orders of algebraic number fields. From the 1990s on, factorization theory expanded its scope into commutative ring theory and since then, multiplicative ideal theory and factorization theory developed in parallel and stimulated each other. At present, both areas are in the midst of a stormy development. This is documented by recent research monographs of Chabert, Elliott, Fontana et al., Geroldinger et al., Grynkiewicz, Halter-Koch, Knebusch et al., Wang and Kim ([103, 104, 53, 21, 38, 82, 88, 139, 63]), by a steady flow of conferences devoted to these fields and the associated conference proceedings (for some of the last 10 years, see [51, 28, 52, 45, 8, 22, 19]).
The above mentioned monographs contain an abundance of open problems (in addition, we would like to refer to the survey [20] by Cahen et al. on unsolved questions). Thus we made no attempt to give a comprehensive overview on open problems, but we tried to offer brief discussions of some research directions which we hope will stimulate further research in multiplicative ideal and factorization theory.
2. Preliminaries
Let $\mathbb{N}$ denote the set of positive integers and set $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For integers $a,b\in\mathbb{Z}$, let $[a,b]=\{x\in\mathbb{Z}\colon a\leq x\leq b\}$ denote the discrete interval between $a$ and $b$. For subsets $A,B$ of an additive abelian group $G$, let $A+B=\{a+b\colon a\in A,b\in B\}\subseteq G$ denote their sumset. If $S$ is any set, then $\mathcal{P}(S)$ denotes the system of all subsets of $S$ and $\mathcal{P}_{\text{\rm fin}}(S)$ the system of all finite, nonempty subsets of $S$.
All semigroups are commutative and have an identity element. If not stated otherwise, we use multiplicative notation. Let $S$ be a semigroup. We denote by $S^{\times}$ the group of invertible elements, and we say that $S$ is reduced if $S^{\times}=\{1_{S}\}$. Then $S_{{\text{\rm red}}}=S/S^{\times}=\{aS^{\times}\colon a\in S\}$ denotes the associated reduced semigroup of $S$. We say that $S$ is cancellative if $a,b,c\in S$ and $ab=ac$ implies that $b=c$. By a monoid, we mean a cancellative semigroup. Let $M$ be a monoid. An element $a\in M$ is called an atom (or irreducible) if $a\notin M^{\times}$ and an equation $a=bc$, with $b,c\in M$, implies that $b\in M^{\times}$ or $c\in M^{\times}$. We denote by $\mathcal{A}(M)$ the set of atoms and by $\mathsf{q}(M)$ the quotient group of $M$.
Next we introduce a key concept of multiplicative ideal theory, namely ideal systems or, equivalently, star-operations (*-operations). The term star-operation was popularized by Gilmer, following the term Strich-operation introduced by Krull. Following Aubert, Halter-Koch used the term ideal system. These concepts were generalized to semistar operations and then to abstract closure operations by Elliott. Ideal systems were generalized to weak ideal systems and then to module systems by Halter-Koch [86, 88]. In the present survey we use both the terms star-operation and ideal system.
Let $M$ be a monoid. An ideal system on $M$ is a map $r\colon\mathcal{P}(M)\to\mathcal{P}(M)$, defined by $X\mapsto X_{r}:=r(X)$ for all $X\subseteq M$, such that the following conditions are satisfied for all subsets $X,Y\subseteq M$ and all $c\in M$:
  • $X\subseteq X_{r}$.
  • $X\subseteq Y_{r}$ implies $X_{r}\subseteq Y_{r}$.
  • $cM\subseteq\{c\}_{r}$.
  • $cX_{r}=(cX)_{r}$.
Let $r$ be an ideal system on $M$. A subset $I\subseteq M$ is called an $r$-ideal if $I_{r}=I$. We denote by $\mathcal{I}_{r}(M)$ the set of all nonempty $r$-ideals, and we define $r$-multiplication by setting $I\cdot_{r}J=(IJ)_{r}$ for all $I,J\in\mathcal{I}_{r}(M)$. Then $\mathcal{I}_{r}(M)$ together with $r$-multiplication is a reduced semigroup with identity element $M$.
A subset $X\subseteq\mathsf{q}(M)$ is called ($M$-)fractional if $cX\subseteq M$ for some $c\in M$. We define
$X_{r}:=c^{-1}(cX)_{r}\,,$
and note that the definition does not depend on the choice of $c$. A subset $I\subseteq\mathsf{q}(M)$ is a fractional $r$-ideal if $I=X_{r}$ for some fractional subset $X\subseteq\mathsf{q}(M)$. We denote by $\mathcal{F}_{r}(M)$ the set of fractional $r$-ideals. For $I,J\in\mathcal{F}_{r}(M)$, the set $IJ$ is $M$-fractional, and we call $I\cdot_{r}J:=(IJ)_{r}\in\mathcal{F}_{r}(M)$ the $r$-product of $I$ and $J$. A fractional $r$-ideal $J\in\mathcal{F}_{r}(M)$ is called $r$-invertible if there exists a fractional $r$-ideal $J^{\prime}$ such that $J\cdot_{r}J^{\prime}=M$. $(\mathcal{F}_{r}(M),\cdot_{r})$ is a semigroup containing $\mathcal{I}_{r}(M)$, $\mathcal{F}_{r}(M)^{\times}$ is the group of $r$-invertible fractional $r$-ideals, and $\mathcal{I}_{r}^{*}(M)=\mathcal{F}_{r}^{\times}(M)\cap\mathcal{I}_{r}(M)$ the monoid of $r$-invertible $r$-ideals of $M$ with $r$-multiplication.
The map $\delta\colon\mathsf{q}(M)\to\mathcal{F}_{r}(M)^{\times}$, defined by $\delta(a)=aM$, is a group homomorphism with kernel $\ker(\delta)=M^{\times}$. Its cokernel
$\mathcal{C}_{r}(M)=\mathcal{F}_{r}(M)^{\times}/\delta(\mathsf{q}(M))$
is called the $r$-class group of $M$. For subsets $X,Y\subseteq\mathsf{q}(M)$, we denote by $X_{s}=XM$, by $(X\negthinspace:\negthinspace Y)=\{x\in\mathsf{q}(M)\colon xY\subseteq X\}$, by $X^{-1}=(M\negthinspace:\negthinspace X)$, by $X_{v}=(X^{-1})^{-1}$ and by $X_{t}=\bigcup_{E\subseteq X,|E|<\infty}E_{v}$. We will use the $s$-system, the $v$-system, the $t$-system, and the $w$-system. The $v$-system was introduced by Krull and $v$-ideals are also called divisorial ideals. The $w$-system was introduced by R. McCasland and F. Wang (see Section 7).
All rings are commutative and have an identity element. For a commutative ring $D$, we denote by $D^{\times}$ its group of units and by $D^{\bullet}$ its monoid of regular elements. Let $D$ be a domain with quotient field $K$. We denote by $\mathcal{A}(D)$ the set of atoms, by $\mathsf{q}(D)$ the quotient field of $D$, by $\overline{D}$ the integral closure of $D$, and by $\widehat{D}=\widehat{D^{\bullet}}\cup\{0\}$ the complete integral closure of $D$.
An ideal system on $D$ is a map $r\colon\mathcal{P}(D)\to\mathcal{P}(D)$ such that the following conditions are satisfied for all subsets $X,Y\subseteq D$ and all $c\in D$:
  • $X\cup\{0\}\subseteq X_{r}$.
  • $X\subseteq Y_{r}$ implies $X_{r}\subseteq Y_{r}$.
  • $cM\subseteq\{c\}_{r}$.
  • $cX_{r}=(cX)_{r}$.
If $r\colon\mathcal{P}(D)\to\mathcal{P}(D)$ is an ideal system on $D$, then $r^{\bullet}\colon\mathcal{P}(D^{\bullet})\to\mathcal{P}(D^{\bullet})$, defined by $r^{\bullet}(X)=r(X)\setminus\{0\}$ for every $X\subseteq D$, is an ideal system on $D^{\bullet}$. Conversely, if $r\colon\mathcal{P}(D^{\bullet})\to\mathcal{P}(D^{\bullet})$ is an ideal system on $D^{\bullet}$, then $r^{\circ}\colon\mathcal{P}(D)\to\mathcal{P}(D)$, defined by $r^{\circ}(X)=r(X)\cup\{0\}$ for every $X\subseteq D$, is an ideal system on $D$. Thus, in the sequel we will not distinguish between $r,r^{\bullet}$, and $r^{\circ}$ and simply speak of $r$-ideals of $D$ and of $D^{\bullet}$. In particular, $\mathcal{I}_{r}(D)$, $\mathcal{F}_{r}(D)$, $\mathcal{I}^{*}_{r}(D)$, and $\mathcal{C}_{r}(D)=\mathcal{C}_{r}(D^{\bullet})$ have their obvious meaning.
The $d$-system $d\colon\mathcal{P}(D)\to\mathcal{P}(D)$, defined by $X_{d}=\langle X\rangle$ for every $X\subseteq D$, is the system of classical ring ideals of $D$. Whenever we speak of ideals of $D$, we mean $d$-ideals and in all notation introduced above we omit the $d$-suffix. In particular, $\mathcal{I}^{*}(D)=\mathcal{I}_{d}^{*}(D)$ is the monoid of invertible ideals of $D$ and $\mathcal{C}_{d}(D)=\operatorname{Pic}(D)$ is the Picard group of $D$.
A monoid (resp. a domain) is said to be
  • Mori (resp. $v$-Noetherian) if it satisfies the ascending chain condition on $v$-ideals.
  • strong Mori (resp. $w$-Noetherian) if it satisfies the ascending chain condition on $w$-ideals.
  • Krull if it is completely integrally closed and Mori.
Every $v$-ideal is a $t$-ideal and if the monoid is Mori, then $v=t$. Krull monoids and Noetherian domains are strong Mori, and every strong Mori monoid is Mori, because every $t$-ideal is a $w$-ideal.
3. Intersecting Valuation Domains
Let $D$ be a domain with quotient field $K$ and suppose that $D\subsetneq K$. In general, there are a lot of domains between $D$ and $K$. For example, if $D$ is a Noetherian domain with Krull dimension greater than one, then there are uncountably many domains between $D$ and $K$.
A valuation domain $V$ with quotient field $L$ is, in a sense, close to being all of $L$. For example, every domain between $V$ and $L$ is again a valuation domain and is a localization of $V$. And if $V$ has finite Krull dimension then there are only finitely many such intermediate domains.
A theorem of Krull should make the study of valuation domains worthwhile.
Theorem 3.1.
Every integrally closed domain $D$ (not a field) can be expressed as the intersection of all the valuation domains between $D$ and its quotient field.
It seems reasonable then that valuation domains should play a central role in the study of integrally closed domains. They are the indivisible building blocks from which everything is made. Some work has been done in this direction but not a lot. A principal difficulty is that given an integrally closed domain $D$ it is usually hard to determine what the valuation domains are. And, given a collection of valuation domains it is hard to discern what the structure of the intersection is. A valuation overring $V$ of $D$ is local, and its maximal ideal contracts to a prime ideal of $D$. But, often, the maximal ideals of uncountably many valuation overrings all contract to the same prime ideal of $D$.
Clearly, many efforts are made to build new examples of domains as overrings of well-understood domains. This is generally done by starting with a domain that is well-understood and adding in some fractions to obtain an overring. Instead of this bottom-up approach we explore here a top-down approach of intersecting valuation domains.
PROBLEM 1.
  1. (1)
    Given an integrally closed domain what can we say about the collection of its valuation overrings?
  2. (2)
    Given a collection of valuation domains what can we say about the domain that results from intersecting them?
These are both difficult questions about which little is known. In this note we focus on one facet of the second one. As we observed above, valuation domains are local. One could then hope that, given a domain, we could obtain valuation overrings by localizing at prime ideals. This rarely succeeds. But sometimes it does. We say that a domain is a Prüfer domain if the localization at any prime ideal is a valuation domain. So, if given a Prüfer domain, it is easy to write the domain as an intersection of valuation domains. The representation is simply the intersection of the localizations at prime ideals. The other direction is a fascinating question.
PROBLEM 2.
Given a collection of valuation domains, what are necessary and sufficient conditions that we can put on the collection that would make the intersection a Prüfer domain?
The following theorem gives partial answers to Problem 2.
Theorem 3.2.
Let $\{V_{i}\ |\ i\in\Omega\}$ be a collection of valuation domains that all have the same quotient field. In each of the following cases the intersection of the valuation domains is a Prüfer domain.
  1. (1)
    $\Omega$ is finite. [114]
  2. (2)
    There is a monic non-constant polynomial $f\in K[X]$, whose coefficients lie in each $V_{i}$ and which has no root in the residue field of any $V_{i}$. [78]
  3. (3)
    There is a field $L$, whose cardinality is strictly larger than the cardinality of $\Omega$, such that $L\subseteq V_{i}$ for each $i\in\Omega$. [115]
The common theme of these three properties (1) - (3) seems to be that the collection of valuation domains needs to be, in some sense, thin or sparse. This observation, in turn, might lead one to seek a topological condition. We digress to describe a construction of Krull which shows that looking at the standard topologies on a collection of valuation domains is not sufficient to determine when the intersection is a Prüfer domain.
We start by giving a quick demonstration that Prüfer domains have properties that make them easy to work with.
Let $K$ be a field and let $D=K[X,Y]$ be the polynomial ring in $X$ and $Y$. Let $J=(X,Y)$ and $I=(X^{2},Y^{2})$ be ideals of $D$. It is easy to see that $J^{3}=JI$. It seems intuitively clear then that $J^{2}=I$ should be true. But it is not. On the other hand, suppose that $T$ is a Prüfer domain and let $d,r\in T$ be nonzero. Let $B=(d,r)$ and $H=(d^{2},r^{2})$ be ideals of $T$. As above, $BH=B^{3}$. is clear. Now, however, in the context of a Prüfer domain, we can cancel $B$ and conclude that $B^{2}=H$ is true.
Krull’s idea was to define certain closure operations on the collection of fractional ideals of a domain. These closure operations are known as star operations. Of particular interest are the star operations he identified as ab (arithmetisch brauchbar = arithmetically useful). Let $D$ be an integrally closed domain and let * be an ab star operation. Then when $I,J,H$ are finitely generated ideals such that $(IH)^{*}\subseteq(JH)^{*}$ then we can conclude that $I^{*}\subseteq J^{*}$.
Before pressing on, we recap what we have discussed so far and where we want to go with it.
  • Every integrally closed domain is the intersection of its valuation overrings.
  • Prüfer domains are precisely the domains for which any localization at a prime ideal is a valuation domain.
  • We would like to know what conditions on a collection of valuation rings with a common quotient field characterize those collections for which the intersection is a Prüfer domain.
  • We have some results which seem to indicate that the property (or properties) that work will be topological in nature.
There are two well-known topologies which can be placed on a collection of valuation domains (for a sample of contributions in this direction see [33, 37, 50, 49, 81]). These are known as
  1. (1)
    the Zariski topology, and
  2. (2)
    the constructible or patch or ultrafilter topology.
Some amazing work of Krull demonstrates clearly that neither of these topologies alone can provide a solution to our problem.
Let $V$ be a valuation domain. For a polynomial $f\in V[X]$, let $c(f)$ be the ideal of $V$ generated by the coefficients of $f$. Let $N=\{f\in V[X]\mid c(f)=V\}$ and finally, let $V(X)=N^{-1}V[X]$. $V(X)$ is known as the Nagata extension of $V$. $V(X)$ is a valuation domain which is almost identical to $V$. The only significant difference between $V$ and $V(X)$ is that $V(X)$ has more units.
Let $D$ be an integrally closed domain (not a field) with quotient field $K$. Krull showed how to use ab star operations on $D$ to build what he called Kronecker function domains. [54] There is always a unique minimal Kronecker function domain which we label as $\operatorname{Kr}(D)$ which has the following properties.
  1. (1)
    $\operatorname{Kr}(D)$ is an overring of $D[X]$.
  2. (2)
    $\operatorname{Kr}(D)\cap K=D$.
  3. (3)
    $\operatorname{Kr}(D)$ is a Prüfer domain.
  4. (4)
    The localizations $\operatorname{Kr}(D)_{\mathfrak{p}}$ where $\mathfrak{p}$ runs over the prime ideals of $\operatorname{Kr}(D)$ are precisely the valuation domains $V(X)$, where $V$ runs over the valuation overrings of $D$.
The key importance of $\operatorname{Kr}(D)$ for us is that when $D$ is an integrally closed domain the space of valuation overrings of $D$ and the space of valuation overrings of $\operatorname{Kr}(D)$ are homeomorphic for both of the topologies noted above. This is problematic because, while $D$ can be any integrally closed domain, $\operatorname{Kr}(D)$ is always a Prüfer domain. Hence, there is no hope of characterizing when an intersection of valuation domains is a Prüfer domain using just information about the topology of the collection of valuation domains.
4. Classification of Integral Domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$
We survey recent developments in the classification of domains that lie strictly between the polynomial ring $\mathbb{Z}[X]$ and its field of fractions $\mathbb{Q}[X]$. We focus on two major studies:
  • [109]
    which classifies integrally closed domains (including Prüfer and Dedekind domains) via order-theoretic arguments, extending classical valuation-theoretic methods.
  • [118]
    which classifies Dedekind domains using the concept of generalized rings of integer-valued polynomials, broadening earlier results on factorization and class groups.
Both works build on classical ideas from valuation theory, factorization theory, and the arithmetic of polynomial rings, offering new insights into how certain local constructions (valuation overrings) glue together to form global integrally closed domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$.
The primary motivation of [109] is to give an order-theoretic classification of integrally closed domains situated between $\mathbb{Z}_{(p)}[X]$ and $\mathbb{Q}[X]$, where $p$ is a prime number and $\mathbb{Z}_{(p)}$ denotes the localization of $\mathbb{Z}$ at $p\mathbb{Z}$. In particular, [109] focuses on the structure of valuation overrings and extends classical results of MacLane on constructing valuations via key polynomials.
A representative theorem from [109] is:
Theorem 4.1 ([109]).
Let $p\in\mathbb{Z}$ be a fixed prime, and let $D$ be a domain such that
$\mathbb{Z}_{(p)}[X]\;\subseteq\;D\;\subseteq\;\mathbb{Q}[X].$
Then $D$ is Prüfer if and only if $D$ is integrally closed and all the $p$-unitary valuation overrings of $D$ are limit.
Constructive aspects hinge on the notion of key polynomials, which allow for systematic building of valuation overrings, ultimately classifying the possible integrally closed domains in a step-by-step manner.
Having explored the order-theoretic classification of integrally closed domains, we now shift our focus to Dedekind domains. To facilitate this discussion, we first introduce the relevant notation.
Let $\mathbb{P}$ be the set of all prime numbers. For a fixed $p\in\mathbb{P}$, we adopt the following notation:
  • $\mathbb{Z}_{p}$ denotes the ring of $p$-adic integers.
  • $\mathbb{Q}_{p}$ denotes the field of $p$-adic integers.
  • $\mathbb{C}_{p}$ denotes the completion of the algebraic closure of the field $\mathbb{Q}_{p}$.
  • ¯$\overline{\mathbb{Z}}_{p}$ denotes the absolute integral closure of $\mathbb{Z}_{p}$.
  • ^¯$\overline{\widehat{\mathbb{Z}}}=\prod_{p\in\mathbb{P}}\overline{\mathbb{Z}}_{p}$.
A generalized ring of integer-valued polynomials is defined, for a subset $\underline{E}\subset\overline{\widehat{\mathbb{Z}}}$, as
$\mathrm{Int}_{\mathbb{Q}}(\underline{E},\overline{\widehat{\mathbb{Z}}})\;=\;\{\,f\in\mathbb{Q}[X]\,\mid\,f(\alpha)\in\overline{\widehat{\mathbb{Z}}}\;\text{for all}\;\alpha\in\underline{E}\}.$
This idea generalizes the classical ring of integer-valued polynomials.
Let $\underline{E}=\prod_{p}E_{p}$ be a subset of $\overline{\widehat{\mathbb{Z}}}$. Following [117, Definition 2.11], we say that $\underline{E}$ is polynomially factorizable if, for every $g\in\mathbb{Z}[X]$ and every $\alpha=(\alpha_{p})\in\underline{E}$, there exist $n,d\in\mathbb{Z}$ with $n,d\geq 1$ such that $g(\alpha)^{n}/d$ is a unit of $\overline{\widehat{\mathbb{Z}}}$.
In [117], the author classifies Dedekind domains with finite residue fields of prime characteristic that lie between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$. These domains are precisely realized as generalized rings of integer-valued polynomials. The work is inspired by earlier studies on factorization theory, integer-valued polynomials, and class group computations.
A key observation is that if $W$ is a one-dimensional valuation overring of such a Dedekind domain with $\mathbb{Q}\subset W$, then $W=\mathbb{Q}[X]_{(q)}$ for some irreducible polynomial $q\in\mathbb{Q}[X]$, implying that its residue field has infinite cardinality. In [118], the author extends this classification by relaxing the condition on the cardinality of residue fields of prime characteristic.
Recall that an element $\alpha\in\mathbb{C}_{p}$ has bounded ramification if the extension $\mathbb{Q}_{p}\subseteq\mathbb{Q}_{p}(\alpha)$ has finite ramification. Denote by $\mathbb{C}_{p}^{br}$ the set of all elements of $\mathbb{C}_{p}$ with bounded ramification, and let $\mathbb{O}_{p}$ be the unique local ring of $\mathbb{C}_{p}$. Define $\mathcal{O}=\prod_{p\in\mathbb{P}}\mathbb{O}_{p}\subset\prod_{p\in\mathbb{P}}\mathbb{C}_{p}$.
Theorem 4.2 ([118, Theorem 3.4]).
Let $R$ be a Dedekind domain such that $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$. Then $R$ is equal to $\operatorname{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ for some polynomially factorizable subset $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}\subset\mathcal{O}$, where for each prime $p$, the set $E_{p}\subset\mathbb{O}_{p}\cap\mathbb{C}_{p}^{br}$ consists of finitely many transcendental elements over $\mathbb{Q}$.
This result completes the classification of Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ and provides a powerful characterization of Dedekind domains in this setting, linking local valuation data and integer-valued polynomial constraints to global structural properties such as the class group. The result is obtained by characterizing discrete valuation rings (DVRs) $W$ of the field $\mathbb{Q}(X)$ that are residually algebraic over $\mathbb{Z}_{(p)}$ for some prime $p$ (see [118, Corollary 2.28]).
When the extension of residue fields has infinite degree, such a domain is given by
$\{\phi\in\mathbb{Q}(X)\mid v_{p}(\phi(\alpha))\geq 0\}$
for some $\alpha\in\mathbb{C}_{p}$, which is transcendental over $\mathbb{Q}_{p}$. The case where the residue field extension has finite degree was addressed in [116].
This result also generalizes findings by Loper, Werner, and others concerning Prüfer domains and valuation-theoretic constructions.
Although [109] and [117, 118] offer significant advances in classifying integrally closed domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$, the broader landscape of such domains is still only partially explored. In particular, unifying generalized integer-valued polynomial viewpoints with valuation-theoretic frameworks remains an active area of research, as do possible applications to:
  • Krull or ((((strong)))) Mori domains: Extending classification results to settings with additional constraints (e.g., finite residue fields, specific $v$-class groups, etc.).
  • Factorial domains: Investigating when generalized integer-valued polynomial rings are factorial.
Recall that an integral domain $D$ is called an almost Dedekind domain (resp., a $t$-almost Dedekind domain) if $D_{\mathfrak{m}}$ is a DVR for all maximal ideals (resp., maximal $t$-ideals) $\mathfrak{m}$ of $D$.
PROBLEM 3.
Classify all integral domains of a specified type (e.g., ($t$-)almost Dedekind domains and Krull domains satisfying additional conditions) and determine their placement between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$. Furthermore, investigate the extent to which such domains can be realized as:
  • Generalized rings of integer-valued polynomials, building on [117, 118], or
  • Valuation overrings constructed via MacLane-type key polynomial techniques, as in [109].
5. Generalizing Kronecker Function Domains
In a previous section we discussed a construction of Krull known as the Kronecker function domain. We start with an integrally closed domain $D$ with quotient field $K$ and build the domain $\operatorname{Kr}(D)$ which satisfies
  1. (1)
    $\operatorname{Kr}(D)$ lies between $D[X]$ and $K(X)$.
  2. (2)
    $\operatorname{Kr}(D)$ is a Prüfer domain.
  3. (3)
    The prime ideals of $\operatorname{Kr}(D)$ correspond precisely with the valuation overrings of $D$.
There are two features of this that might strike one as annoyances that one would wish to avoid. We start as above with an integrally closed domain $D$ with quotient field $K$.
  • The actual construction is built using the ab star operations. It is cumbersome. One could wish for a simpler construction.
  • One could wonder whether the base ring $D$ is essential. Is it possible to build a Kronecker function domain using just $K$ without $D$?
In 2003 Halter-Koch [85] gave what he called an axiomatic treatment of what he called Kronecker domains which gives us the construction we were looking for above.
Let $K$ be a field which has valuation subrings. Let $X$ be an indeterminate. We say that a ring $T$ is a Kronecker domain provided
  1. (1)
    $T\subseteq K(X)$.
  2. (2)
    $X$ is a unit in $T$.
  3. (3)
    For every nonzero polynomial $f\in K[X]$, we have $f(0)\in fT$.
If these conditions are satisfied, then we have also
  • $T$ is a Prüfer domain.
  • If $\mathfrak{p}$ is a prime ideal of $T$ then $T_{\mathfrak{p}}$ has the form $V(X)$ for some valuation domain $V$ contained in $K$.
  • The conventional Kronecker function domains defined for subrings $D$ of $K$ are included as Kronecker domains.
Halter-Koch’s notion of Kronecker domain has been known since 2003 and very little has been done in that time to study them. Actually, Halter-Koch himself extended what is known by himself by demonstrating that they can be built with a variant of the conventional construction ([88, Theorem 6.5.5]). Many questions could be posed. We will content ourselves with one.
PROBLEM 4.
Fabbri and Heubo-Kwegna [41] give explicit details for the construction of a non-traditional Kronecker domain in a very specialized setting. Can their results be generalized to a much wider setting?
6. Coherent Domains
Coherent domains were first defined by Chase [30] in 1960. A domain $D$ is coherent if it satisfies one of the following three equivalent conditions (note there are many more equivalent ones).
  1. (a)
    The intersection of any two finitely generated ideals of $D$ is finitely generated.
  2. (b)
    Any finitely generated ideal of $D$ is finitely presented.
  3. (c)
    The direct product of any family of flat $D$-modules is flat.
We focus on the first of these conditions. It is clear from this condition that any Noetherian domain is coherent. It is not as simple, but still straightforward to prove that any Prüfer domain is coherent. When one is considering all domains there are other natural examples of coherent domains. In the realm of one-dimensional domains things are much more restrictive. A one-dimensional coherent domain which is integrally closed is a Prüfer domain. The integral closure of a one-dimensional Noetherian domain is a Dedekind domain which is both Noetherian and Prüfer. The class of one-dimensional coherent domains which are neither Noetherian nor Prüfer is vanishingly small. We are aware of one such example in the literature. This example is not integrally closed but has Prüfer integral closure.
This state of affairs led Vasconcelos to ask
PROBLEM 5.
Is the integral closure of a one-dimensional coherent domain a Prüfer domain? [79]
One might hope to make progress on this question by considering many examples of one-dimensional coherent domains that are neither integrally closed nor Noetherian. But again, such examples scarcely exist.
Instead of looking at this question from a forward direction (examine the integral closure of a one-dimensional coherent domain) maybe one can make progress by looking at the question from a backwards perspective.
  • Let $D$ be a one-dimensional integrally closed domain which is not Prüfer. Is $D$ in any sense close to being coherent?
A recent result shows that if such a domain $D$ is local it is, in a sense anti-coherent.
Theorem 6.1.
Let $D$ be a one-dimensional integrally closed local domain which is not a Prüfer domain. Let $I,J$ be two principal ideals of $D$ with no containment relations. Then $I\cap J$ is not finitely generated. [83]
This seems to be evidence toward a positive answer to Vasconcelos’ question. How could a coherent domain - where the intersection of two finitely generated ideals is always finitely generated - have as its integral closure a domain where the intersection of two principal ideals is almost never finitely generated?
PROBLEM 6.
Let $K$ be a field. Let $X,Y$ be indeterminates. Consider the valuation domain $V=K(X)[Y]_{(Y)}$. Let $M$ be the maximal ideal of $V$. Let $D=K+M$. Then $D$ is a one-dimensional local domain which is integrally closed and not Prüfer. Can we prove that $D$ is not the integral closure of a coherent domain?
7. Theory of $w$-modules
Originally formulated by McCasland and Wang [110], [111], $w$-module theory refines classical module and ideal theories in commutative algebra by focusing on a specific star operation, the $w$-operation. The concept of $w$-ideals coincides with the $F_{\infty}$-operation previously considered by J.R. Hedstrom and E.G. Houston [90], and the $w$-ideals coincide with the semidivisorial ideals of S. Glaz and W. Vasconcelos [80]).
The framework of $w$-ideals generalizes earlier results on $t$-ideals and divisorial ideals while providing new insights in non-Noetherian settings. The core motivation for $w$-module theory arose from studying the structure of commutative rings and the interplay of their star operations with modules and ideals.
Definition 7.1.
Let $R$ be a commutative ring, possibly with zero divisors.
  1. (1)
    A nonzero ideal $J$ of $R$ is called a Glaz–Vasconcelos ideal (GV-ideal) if $J$ is finitely generated and the natural homomorphism
    $\varphi\colon R\;\longrightarrow\;\operatorname{Hom}_{R}(J,R)$
    is an isomorphism. The set $\operatorname{GV}(R)$ of all GV-ideals forms a multiplicative system of ideals in $R$.
  2. (2)
    An $R$-module $M$ is called GV-torsion-free if, for any $J\in\operatorname{GV}(R)$, the condition $Jx=0$ (for some $x\in M$) implies $x=0$. On the other hand, $M$ is called GV-torsion if, for every nonzero element $x\in M$, there exists $J\in\operatorname{GV}(R)$ such that $Jx=0$. It is evident that the pair (GV-torsion modules, GV-torsion-free modules) defines a hereditary torsion theory, first introduced in [102, p. 1618].
  3. (3)
    For a GV-torsion-free $R$-module $M$, define
    $M_{w}\;=\;\bigl{\{}\,x\in E(M)\;\bigm{|}\;\text{there exists }J\in\operatorname{GV}(R)\text{ such that }Jx\subseteq M\bigr{\}},$
    where $E(M)$ is the injective envelope of $M$. The module $M_{w}$ is called the $w$-envelope of $M$.
  4. (4)
    A GV-torsion-free $R$-module $M$ is called a $w$-module if $\operatorname{Ext}_{R}^{1}(R/J,M)=0$ for every $J\in\operatorname{GV}(R)$; equivalently, if $M=M_{w}$. In particular, an ideal $I$ of $R$ is called a $w$-ideal if $I=I_{w}$.
  5. (5)
    An $R$-module $M$ is said to be $w$-flat if, for every maximal $w$-ideal $\mathfrak{m}$ of $R$, the localization $M_{\mathfrak{m}}$ is flat over $R_{\mathfrak{m}}$.
The $w$-operation builds a bridge towards homological algebra. The resulting $w$-module theory can be used to study certain “high-dimensional” domains from a module-theoretic perspective. In fact, the $w$-module theory is a type of locally homological method.
Every flat $R$-module is $w$-flat, and if $D$ is a Prüfer domain, $w$-flat $D$-modules coincide with flat $D$-modules. This shows that $w$-module theory specializes to classical notions in more restricted settings.
It is well-known that Noetherian rings play an important role in commutative algebra and algebraic geometry. Using $w$-operations, strong Mori domains were introduced in [140]. It was previously claimed ([57, Corollary IX.5.5]) that for a domain $D$ with quotient field $K$, the injective module $E(K/D)$ is $\Sigma$-injective if and only if $D$ is a Mori domain. However, this statement is incorrect: one must replace ‘Mori’ with ‘strong Mori’ (see [56, Theorem 9]). In recent years, many classical theorems about Noetherian rings have been generalized to strong Mori domains.
In parallel with $w$-module theory, the study of Krull domains has been pivotal in commutative algebra via the lens of star operations. Indeed, a Krull domain $D$ can be characterized by the following conditions.
  1. (i)
    For every height-one prime ideal $\mathfrak{p}$, the localization $D_{\mathfrak{p}}$ is a discrete valuation domain.
  2. (ii)
    $D=\bigcap D_{\mathfrak{p}}$, the intersection taken over all height-one prime ideals $\mathfrak{p}$.
  3. (iii)
    Every nonzero non-unit in $D$ is contained in only finitely many height-one prime ideals.
Analogous to how Dedekind domains can be characterized by the invertibility of nonzero ideals, Krull domains can be characterized as follows ([101, Theorem 3.6], [139, Theorem 7.9.3]):
[…]
This implies that a Krull domain can essentially be considered a $t$- (or $w$-) Dedekind domain. Furthermore, it suggests that Krull domains can be studied in a manner similar to Dedekind domains, utilizing star-operator methods.
PROBLEM 7.
Prove or disprove the assertion that
  • Every result in classical commutative algebra has a corresponding analogue in $w$-module theory.
Finding an accurate analogue for this may be challenging and could require the development of new concepts and techniques. For instance, in a Dedekind domain, any finitely generated module $M$ decomposes as a direct sum of a projective module and a torsion module. A central question in $w$-module theory (and more broadly in star-operation settings) is whether such classical decompositions have analogues in Krull domains. For Krull domains, what is the corresponding formulation for finitely generated ($w$-finitely generated) modules?
Recent extensions explore $w$-factor rings. For a domain $D$, the following conditions are equivalent.
  1. (a)
    $D$ is Krull.
  2. (b)
    The $w$-factor ring $(D/I)_{w}$ is a zero-dimensional semilocal principal ideal ring for every proper $w$-ideal $I$ of $D$, [26, Theorem 4.5].
  3. (c)
    $(D/I)_{w}$ is a quasi-Frobenius ring [148, Theorem 9].
[…]
PROBLEM 8.
[…]
8. Homological versus Elementary Methods
Every ring considered in this section is an integral domain.
In 1937 Krull defined a regular local ring to be a Noetherian local ring $D$ with maximal ideal $\mathfrak{m}$ for which the minimum number of elements in a generating set for $\mathfrak{m}$ is the same as the height of $\mathfrak{m}$ [106]. This definition is completely elementary.
In 1956 Serre [133] proved that a Noetherian local ring is regular if and only if it has finite global dimension. This homological characterization of regular local rings was pursued because it led to an easy proof that a localization of a regular local ring at a prime ideal was still regular, which had proven to be quite elusive using non-homological methods.
This event is often cited as validating the use of homological methods in the study of ring theory. In fact, homological methods have become quite dominant. We propose several settings in which the elementary methods might regain prominence.
Question I. Does there now exist a non-homological proof that any localization of a regular local ring is still regular?
Answer: There exist informal sites on the internet that indicate that the answer is yes. These sites indicate that the proof involves careful use of regular sequences. This is not surprising since it is well-known that an ideal generated by a regular sequence has finite projective dimension. But we were unable to find a published proof. So
PROBLEM 9.
Can one prove non homologically that any localization of a regular local ring is still regular?
We push this a little farther. Ratliff [126] defined a set $\{d_{0},d_{1},\ldots,d_{n}\}$ of elements in a local domain $(D,\mathfrak{m})$ to be analytically independent if, whenever $F(d_{0},\ldots,d_{n})=0$ for some homogeneous polynomial $F(X_{0},\ldots,X_{n})\in D[X_{0},\ldots,X_{n}]$, then all coefficients of $F$ are in $\mathfrak{m}$. This generalizes the notion of a regular sequence. Say that the analytically independent dimension is the maximal size of an analytically independent set. Is it possible that the notion of finite projective dimension could be replaced by the notion of analytically independent dimension?
PROBLEM 10.
An ideal generated by a regular sequence has finite projective dimension. A regular sequence is analytically independent. Is it true that an ideal generated by an analytically independent set has finite projective dimension?
Question II: Can we use homological terminology to define classes of rings that are not being studied at all and might be profitably studied with elementary techniques?
Answer: Perhaps. Recall that a Dedekind domain is a domain for which every nonzero ideal is invertible. The more general notion of a Prüfer domain is a domain for which every finitely generated ideal is invertible. We could similarly say that a domain has pseudo-finite global dimension provided there is an upper bound on the projective dimensions of all finitely generated modules. Perhaps the union of an ascending tower of regular local rings could provide nontrivial examples of such rings. Examples of such towers are given in recent papers concerning the so-called Shannon extensions [91].
PROBLEM 11.
Can the property of having pseudo-finite global dimension be defined in elementary terms?
PROBLEM 12.
Do Shannon extensions have pseudo-finite global dimension?
9. Gorenstein Multiplicative Ideal Theory
Gorenstein homological algebra arose as an extension of classical homological algebra to better accommodate non-regular contexts. Traditional projective, injective, and flat modules often fail to capture subtler behaviors in more general rings and modules, especially those beyond Noetherian or regular situations. Gorenstein dimensions (and related Gorenstein notions) were introduced to address these shortcomings, building on Auslander and Bridger’s foundational work on Gorenstein projective modules ([7]). Their investigations were initially guided by syzygy theory over Noetherian rings and the concept of complete intersections.
Subsequent expansions by Enochs, Jenda, and Torrecillas introduced a range of new tools—such as relative homological dimensions ([39, 40]). Classical homological algebra heavily emphasizes projective, injective, and flat modules. The Gorenstein versions of these notions ‘relax’ certain regularity assumptions.
Throughout this section,
let $R$ denote a commutative ring, possibly with zero divisors.
Definition 9.1.
Let $M$ be an $R$-module.
  1. (1)
    $M$ is said to be Gorenstein projective if there exists an exact sequence of projective $R$-modules
    $\cdots\;\longrightarrow\;P_{1}\;\longrightarrow\;P_{0}\;\longrightarrow\;P^{0}\;\longrightarrow\;P^{1}\;\longrightarrow\;\cdots$
    such that $M\cong\ker(P^{0}\to P^{1})$ and $\operatorname{Hom}_{R}(-,Q)$ preserves exactness of this sequence for every projective $R$-module $Q$.
  2. (2)
    $M$ is said to be Gorenstein injective if there exists an exact sequence of injective $R$-modules
    $\cdots\;\longrightarrow\;I_{1}\;\longrightarrow\;I_{0}\;\longrightarrow\;I^{0}\;\longrightarrow\;I^{1}\;\longrightarrow\;\cdots$
    such that $M\cong\ker(I^{0}\to I^{1})$ and $\operatorname{Hom}_{R}(Q,-)$ preserves exactness of this sequence for every injective $R$-module $Q$.
  3. (3)
    $M$ is said to be Gorenstein flat if there exists an exact sequence of flat $R$-modules
    $\cdots\;\longrightarrow\;F_{1}\;\longrightarrow\;F_{0}\;\longrightarrow\;F^{0}\;\longrightarrow\;F^{1}\;\longrightarrow\;\cdots$
    such that $M\cong\ker(F^{0}\to F^{1})$ and exactness is preserved under the functor $-\otimes_{R}E$ for every injective $R$-module $E$.
  4. (4)
    The Gorenstein projective dimension of $M$, denoted $\operatorname{Gpd}_{R}(M)$, is defined as the smallest integer $n$ (if such an $n$ exists) such that there exists an exact sequence of Gorenstein projective modules:
    $0\longrightarrow G_{n}\longrightarrow G_{n-1}\longrightarrow\cdots\longrightarrow G_{1}\longrightarrow G_{0}\longrightarrow M\longrightarrow 0,$
    where each $G_{i}$ is a Gorenstein projective module. If no such finite $n$ exists, we set $\operatorname{Gpd}_{R}(M)=\infty$.
Example 9.2.
  1. (1)
    If $R$ is a Gorenstein ring, then every finitely generated $R$-module has finite Gorenstein projective dimension. Over a regular local ring, the Gorenstein projective dimension agrees with the usual projective dimension.
  2. (2)
    Let $R=K[X]/(X^{2})$ for a field $K$. This is a Frobenius algebra, so $R$ is both Gorenstein projective and Gorenstein injective as an $R$-module.
Definition 9.3.
An integral domain $D$ is called a
  • Gorenstein Dedekind domain (briefly, G-Dedekind domain) if every ideal of $D$ is G-projective.
  • Gorenstein Prüfer domain (briefly, G-Prüfer domain) if $D$ is coherent and every submodule of a flat $D$-module is Gorenstein flat.
The notion of a Gorenstein Krull domain (G-Krull domain for short) is also introduced in [123]. Examples of G-Prüfer and G-Krull domains are provided in [144].
Holm’s metatheorem [92] proclaims that
  • Every result in classical homological algebra has a Gorenstein counterpart.
This sweeping statement has been largely corroborated by numerous analogies and parallels between classical and Gorenstein homological results. Yet there remain instances in which Gorenstein analogies are only partially true or require stricter hypotheses.
For example, in the classical setting, a domain $D$ is Prüfer if and only if it is locally a valuation domain ([139, Corollary 3.7.22]). In the Gorenstein setting, while every G-Prüfer domain is locally G-Prüfer [145, Corollary 11], the converse fails in general [147, Example 4.1]. Hence, the direct classical analogy does not hold perfectly in Gorenstein contexts.
Classically, $D$ is Prüfer if and only if every overring of $D$ is flat [107, Theorem 6.10]. Analogously, it has been shown that if $R$ is G-Prüfer, then every overring of $R$ is Gorenstein flat [145, Theorem 6]. However, it remains open whether the converse also holds. Namely, is a domain $D$ G-Prüfer if every overring of $D$ is Gorenstein flat? This is a natural Gorenstein-homological parallel to a well-known Prüfer criterion, but it awaits resolution.
A key feature of Dedekind domains $D$ is that every nonzero ideal is invertible, which is equivalent to demanding that $D/I$ be quasi-Frobenius for each nonzero ideal $I$ ([148, Theorem 13]). Similarly, in the Gorenstein Dedekind setting, one has:
  • An integral domain $D$ is G-Dedekind if and only if, for every nonzero element $a\in D$, the factor ring $D/(a)$ is quasi-Frobenius ([139, Theorem 11.7.7]). This implies that every nonzero ideal of $D$ is a $v$-ideal. In fact, $D$ is a G-Dedekind domain if and only if it is a Noetherian divisorial domain.
  • In general, a G-Dedekind domain is one-dimensional and Noetherian ([139, Theorem 11.7.7 and Corollary 11.7.8]), yet can exhibit different behaviors regarding integral closure, reminiscent of how Gorenstein rings relax some classical regularity or integrality conditions.
Recent connections show that factor rings of the form $(D/xD)_{w}$ being quasi-Frobenius for all nonzero $x$ in $D$ characterizes G-Krull domains [146, Theorem 2.7]. Such Gorenstein-enabled statements suggest deep analogies between multiplicative ideal theory and Gorenstein homological algebra, but also underscore certain exceptions where classical and Gorenstein theories diverge.
PROBLEM 13.
Develop explicit characterizations of Gorenstein homological properties—with an emphasis on G-Dedekind and G-Prüfer domains—through both ideal-theoretic and ring-theoretic lenses. In particular:
  • Clarify how the $v$-operation and $\ast$-invertibility interact in the Gorenstein setting, and determine whether the $v$-operation fully captures Gorenstein analogues of classical invertibility.
  • Investigate if and how integral closure behaves differently for G-Dedekind domains compared to classical Dedekind domains.
  • Resolve open questions about overring flatness: under what conditions does “every overring is Gorenstein flat” imply that the base domain is G-Prüfer?
  • Investigate whether $\overline{D}$ is a Prüfer domain when $D$ is G-Prüfer.
10. Cotorsion Theory
Cotorsion theory is a branch of homological algebra that uses the $\operatorname{Ext}$ functor to organize module categories (or abelian categories) into two complementary classes of objects, generalizing concepts such as injectivity, flatness, and their relative versions. By drawing on ideas from torsion theories and projective/injective dimensions, cotorsion theory provides a powerful framework for decomposing modules and understanding the structure of extension groups.
Throughout this section, let $R$ denote a commutative ring, possibly with zero divisors.
Definition 10.1.
A cotorsion pair $(\mathcal{A},\mathcal{B})$ in the category of $R$-modules (or in any abelian category) consists of two classes of modules $\mathcal{A}$ and $\mathcal{B}$ such that:
  1. (1)
    $\mathcal{A}\;=\;\bigl{\{}\,M\;\bigm{|}\;\operatorname{Ext}^{1}_{R}(M,B)=0\text{ for all }B\in\mathcal{B}\bigr{\}}$
    and
    =$\mathcal{B}\;=\;\bigl{\{}\,N\;\bigm{|}\;\operatorname{Ext}^{1}_{R}(A,N)=0\text{ for all }A\in\mathcal{A}\bigr{\}}.$
  2. (2)
    • $\mathcal{A}$ is closed under direct sums, summands, and extensions.
    • $\mathcal{B}$ is closed under direct products, summands, and extensions.
The classes $\mathcal{A}$ and $\mathcal{B}$ are often called orthogonal complements of each other with respect to $\operatorname{Ext}^{1}_{R}(-,-)$.
Example 10.2.
  1. (1)
    Let $\mathcal{M}$ be the category of all $R$-modules, $\mathscr{P}$ the class of all projective modules, and $\mathscr{I}$ the class of all injective modules. Then
    $(\mathscr{P},\mathcal{M})\quad\text{and}\quad(\mathcal{M},\mathscr{I})$
    are both cotorsion pairs in $\mathcal{M}$.
  2. (2)
    Over a principal ideal domain $D$, the cotorsion pair $(\mathcal{A},\mathcal{B})=(\mathrm{Flat},\mathrm{Cotorsion})$ satisfies:
    $\mathcal{A}\;=\;\{\text{flat }D\text{-modules}\},\quad\mathcal{B}\;=\;\bigl{\{}\,N\;\bigm{|}\;\operatorname{Ext}^{1}_{R}(F,N)=0\text{ for all flat }F\bigr{\}}.$
Cotorsion pairs underlie many useful constructions in homological algebra, including relative homological dimensions, abelian model categories, and exact categories. By focusing on how $\operatorname{Ext}^{1}_{R}(-,-)$ vanishes, one can obtain powerful decomposition theorems for modules.
A variant of cotorsion theory, called Tor-cotorsion theory, shifts attention from $\operatorname{Ext}$ to $\operatorname{Tor}$ and hence emphasizes torsion properties and tensor products.
Definition 10.3.
An $R$-module $M$ is called Tor-cotorsion if
$\operatorname{Tor}_{1}^{R}(F,M)\;=\;0\quad\text{for every flat $R$-module }F.$
In classical cotorsion theory, one requires $\operatorname{Ext}^{1}_{R}(F,M)=0$ for every flat $R$-module $F$. In Tor-cotorsion theory, one instead demands that $\operatorname{Tor}_{1}^{R}(F,M)=0$ for every flat $R$-module $F$. Both conditions provide different perspectives on how flat modules (or certain torsion-free classes) might interact with a given module, illuminating complementary aspects of homological dimension and torsion phenomena.
Example 10.4.
Let $\mathscr{F}$ be the class of all flat $R$-modules. Then $(\mathscr{F},\mathcal{M})$ is a (classical) Tor-cotorsion pair in the category of $R$-modules. In a Tor-cotorsion context, we focus on $R$-modules $M$ with $\operatorname{Tor}_{1}^{R}(F,M)=0$ rather than on $\operatorname{Ext}^{1}_{R}(F,M)=0$.
Cotorsion theory (and its Tor-cotorsion counterpart) can effectively describe structural aspects of rings that are central in multiplicative ideal theory. For instance:
  1. (1)
    An integral domain $D$ is Dedekind if and only if the global dimension of $D$ is at most $1$. Equivalently, every nonzero ideal of $D$ is projective (i.e., invertible).
  2. (2)
    An integral domain $D$ is Krull if and only if the global weak $w$-projective dimension $w.w.gl.\dim(D)$ is at most $1$; see [141, Theorem 4.3].
These homological characterizations blend naturally with cotorsion-theoretic viewpoints: the condition $gl.\dim(R)\leq 1$ can be recast via orthogonality requirements on $R$-modules, while the $w$-projective setting relies on extension functors in the $w$-module theory sense. Indeed, building robust (Tor-)cotorsion frameworks broadens the scope of such equivalences, reinforcing links between ring-theoretic properties (e.g., invertibility of ideals) and vanishing results in homological algebra.
Although there are many ways to characterize Prüfer domains, cotorsion viewpoints yield elegant homological statements. First, we recall additional module-theoretic definitions:
Definition 10.5.
Let $D$ be a domain with quotient field $K$, and let $M$ be an $D$-module.
  1. (1)
    $M$ is called a cotorsion module if
    $\operatorname{Ext}^{1}_{D}(F,M)\;=\;0\quad\text{for every flat $D$-module }F.$
  2. (2)
    $M$ is called a Warfield cotorsion module if
    $\operatorname{Tor}_{1}^{D}(M,F)\;=\;0\quad\text{for every torsion-free $D$-module }F.$
  3. (3)
    $M$ is called a Lee cotorsion module if
[…]
Theorem 10.6 ([139], Theorem 12.9.12).
For a domain $D$, the following conditions are equivalent.
  1. (a)
    $D$ is a Prüfer domain.
  2. (b)
    Every cotorsion $D$-module is Warfield cotorsion.
  3. (c)
    Every divisible $D$-module is FP-injective.
  4. (d)
    Every $h$-divisible $D$-module is FP-injective.
  5. (e)
    Every Lee cotorsion $D$-module is FP-injective.
Thus, a domain is Prüfer precisely when several natural classes of modules (cotorsion, divisible, etc.) coincide homologically or admit uniform homological dimensions. This result interlinks classical ideals-based characterizations (e.g., localizing at height-one primes or checking invertibility of ideals) with module-theoretic vanishing conditions.
[…]
PROBLEM 14.
Develop explicit cotorsion and Tor-cotorsion characterizations of rings and integral domains that arise in multiplicative ideal theory. In particular, investigate how homological invariants such as
[…]
can be used to classify such rings. Examples of interest include:
  • Prüfer domains: Extend Theorem 10.6 by identifying further cotorsion-theoretic criteria.
  • Krull domains: Relate the $w$-projective dimension viewpoint to
    (Tor-)cotorsion approaches.
[…]
11. Factorization Theory in Mori Monoids
Factorization theory has spread into various subfields of algebra. However, in this survey we restrict ourselves to factorization theory in Mori and Krull monoids and domains. We discuss neither factorization theory in non-commutative ring theory [11, 136, 16], nor combinatorial factorization theory with its overlap to additive combinatorics, [63], nor factorization theory in non-BF-monoids [35], nor factorizations into distinguished building blocks other than irreducibles [32].
We start with a lemma highlighting the abundance of Mori and Krull monoids in commutative ring theory. For more on the ideal theory of Mori monoids we refer to [88, Chapter 5] and for more of this flavor in the setting of $v$-Marot and $t$-Marot rings see [27, 14]. For Mori monoids with a combinatorial flavor see [42]. Moreover, we refer to a survey of Facchini [43] for the role of Krull monoids in module theory.
skip
Lemma 11.1.
  1. (1)
    A domain $D$ is Mori if and only if $D^{\bullet}$ is a Mori monoid if and only if $\mathcal{I}_{v}^{*}(D)$ is a Mori monoid.
  2. (2)
    If $D$ is a $v$-Marot ring, then $D$ is a Mori ring if and only if $D^{\bullet}$ is a Mori monoid.
  3. (3)
    Let $D$ be a domain and let $r$ be an ideal system on $D$ with $\mathcal{I}_{r}(D)\subseteq\mathcal{I}(D)$. Then $D$ is a Krull domain if and only if $D^{\bullet}$ is a Krull monoid if and only if $\mathcal{I}_{r}^{*}(D)$ is a Krull monoid.
  4. (4)
    If $D$ is a Noetherian domain, then its complete integral closure $\widehat{D}$ is Krull and $\mathcal{I}^{*}(D)$ is a Mori monoid.
  5. (5)
    Let $M$ be a Mori monoid. If $M$ is seminormal or if the conductor $(M\colon\widehat{M})$ is nonempty, then the complete integral closure $\widehat{M}$ is Krull. Furthermore, the complete integral closure of a strong Mori domain is Krull.
Proof.
For (1) and (2) we refer to [70] and for (3) to [69, Theorem 4.5]. The first claim in (4) is the Theorem of Mori-Nagata and the second claim follows from [65, Example 2.1].
(5) For a seminormal $M$ the claim follows from [88, Theorem 5.2.8] and if the conductor is nonempty, then it follows from [64, Theorem 2.3.5]. If $M$ is a strong Mori domain, then the claim follows from [139, Theorems 7.7.14 and 7.10.21]. ∎
We start the discussion of Arithmetic Finiteness Properties. Let $M$ be a monoid. We say that $M$ is
  1. (1)
    atomic if every element $a\in M\setminus M^{\times}$ has a factorization into atoms.
  2. (2)
    factorial if one of the following two equivalent conditions hold.
    • (a)
      $M$ is atomic and factorization into atoms is unique up to the order and up to associates.
    • (b)
      $M$ is Krull and each divisorial ideal of $M$ is principal.
The uniqueness statement in Condition 2.(a) means that whenever $a=u_{1}\ldots u_{k}=v_{1}\ldots v_{\ell}$, with $k,\ell\in\mathbb{N}$ and $u_{1},\ldots,u_{k},v_{1},\ldots,v_{\ell}\in\mathcal{A}(M)$, then $k=\ell$ and there is a permutation $\sigma\colon[1,k]\to[1,k]$ such that $u_{i}M^{\times}=v_{\sigma(i)}M^{\times}$ for all $i\in[1,k]$.
Mori monoids are atomic. Indeed, suppose that $M$ is Mori and assume to the contrary that $\Omega=\{aM\colon a\in M\setminus M^{\times}\ \text{is not a product of atoms}\}$ is nonempty. Since principal ideals are divisorial, $\Omega$ has a maximal element $aM\in\Omega$. Then there are $b,c\in M\setminus M^{\times}$ such that $a=bc$, $aM\subsetneq bM$, and $aM\subsetneq cM$. Thus, $b$ and $c$ are products of atoms and so is $a=bc$, a contradiction. The overall goal is to describe the non-uniqueness of factorizations in Mori monoids which are either not Krull or Krull with nontrivial $v$-class group. This is done by arithmetic invariants which are set in context with algebraic parameters of the respective class of Mori monoids.
We introduce length sets and some parameters controlling their structure. If $a=u_{1}\ldots u_{k}$, with $k\in\mathbb{N}$ and $u_{1},\ldots,u_{k}\in\mathcal{A}(M)$, then the number of factors $k$ is called a factorization length of $a$ and the set $\mathsf{L}_{M}(a)=\mathsf{L}(a)$ of all factorization lengths of $a$ is called the length set of $a$. Note that $1\in\mathsf{L}(a)$ if and only if $a$ is an atom if and only if $\mathsf{L}(a)=\{1\}$. It is a convenient convention to set $\mathsf{L}(a)=\{0\}$ if $a$ is invertible. We denote by
$\mathcal{L}(M)=\{\mathsf{L}(a)\colon a\in M\}$
the system of length sets of $M$. The monoid $M$ is said to be half-factorial if $|\mathsf{L}(a)|=1$ for all $a\in M$. Suppose that $M$ is not half-factorial. Then there exists some $a\in M$ with factorizations of two different lengths, say
$a=u_{1}\ldots u_{k}=v_{1}\ldots v_{\ell}\,,$
where $k,\ell\in\mathbb{N}$ with $k<\ell$ and atoms $u_{1},\ldots,u_{k},v_{1},\ldots,v_{\ell}$. Then, for every $N\in\mathbb{N}$, we have
$a^{N}=(u_{1}\ldots u_{k})^{i}(v_{1}\ldots v_{\ell})^{N-i}\quad\text{for all $i\in[0,N]$}\,,$
whence $|\mathsf{L}(a^{N})|>N$. Thus, length sets may become arbitrarily large. For an element $a\in M$, let $\omega_{M}(a)=\omega(a)$ be the smallest $N\in\mathbb{N}_{0}\cup\{\infty\}$ for which the following holds :
  • If $m\in\mathbb{N}$ and $v_{1},\ldots,v_{m}\in M$ with $a\mid v_{1}\ldots v_{m}$, then there exists a subset $\Omega\subseteq[1,m]$ such that $|\Omega|\leq N$ and $a\mid\prod_{\nu\in\Omega}v_{\nu}$.
For an atom $a\in M$, let $\mathsf{t}(a)$ be the smallest $N\in\mathbb{N}_{0}\cup\{\infty\}$ for which the following holds :
  • If $m\in\mathbb{N}$ and $v_{1},\dots,v_{m}\in\mathcal{A}(M)$ are such that  $a\mid v_{1}\ldots v_{m}$, but $a$ divides no proper subproduct of $v_{1}\ldots v_{m}$, then there exist $\ell\in\mathbb{N}$ and $u_{2},\ldots,u_{\ell}\in\mathcal{A}(M)$ such that $v_{1}\ldots v_{m}=au_{2}\cdot\ldots\cdot u_{\ell}$ and $\max\{\ell,\,m\}\leq N$.
We say that $M$ is locally tame if $\mathsf{t}(a)<\infty$ for all $a\in\mathcal{A}(M)$. By definition, $\omega(a)=1$ if and only if $a$ is a prime element. If $a$ is an atom but not prime, then $\omega(a)\leq\mathsf{t}(a)$.
The next lemma gathers some basic arithmetic properties of Mori monoids.
Lemma 11.2.
Let $M$ be a Mori monoid.
  1. (1)
    All length sets are finite and nonempty (((([64, Theorem 2.2.9])))).
  2. (2)
    $\omega(a)<\infty$ for all $a\in M$ (((([65, Theorem 4.2])))).
For a finite, nonempty set $L=\{m_{1},\ldots,m_{k}\}\subset\mathbb{N}$ with $k\in\mathbb{N}$ and $m_{1}<\ldots<m_{k}$, we call
  • $\Delta(L)=\{m_{i+1}-m_{i}\colon i\in[1,k-1]\}\subseteq\mathbb{N}$ the set of distances of $L$, and
  • $\rho(L)=m_{k}/m_{1}$ the elasticity of $L$.
For convenience, we set $\rho(\{0\})=1$. Let $M$ be a Mori monoid and let $k\in\mathbb{N}$. Then
  • $\Delta(M)=\bigcup_{L\in\mathcal{L}(H)}\Delta(L)\ \subseteq\mathbb{N}$ is the set of distances of $M$,
  • $\rho(M)=\sup\{\rho(L)\colon L\in\mathcal{L}(H)\}\in\mathbb{R}_{\geq 1}\cup\{\infty\}$ is the elasticity of $M$, and
  • $\mathcal{U}_{k}(M)=\bigcup_{k\in L,L\in\mathcal{L}(M)}L\ \subseteq\mathbb{N}$ is the union of length sets containing $k$.
Thus, by definition,
$M\ \text{is half-factorial}\quad\text{if and only if}\quad\Delta(M)=\emptyset\quad\text{if and only if}\quad\rho(M)=1\,,$
and, by a straightforward calculation, we get
$\rho(M)=\lim_{k\to\infty}\frac{\sup\mathcal{U}_{k}(M)}{k}\,.$
We say that $M$ satisfies the Structure Theorem for Unions if there are $k^{*}\geq 1$ and $N\geq 0$ such that, for all $k\geq k^{*}$,  $\mathcal{U}_{k}(M)$ has the form
$\mathcal{U}_{k}(M)=y_{k}+\big{(}L^{\prime}\cup L^{*}\cup L^{\prime\prime}\big{)}\subset y_{k}+d\mathbb{Z},$
where $y_{k}\in\mathbb{Z}$, $L^{*}$ is a nonempty (finite or infinite) arithmetic progression with difference $d$,  $\min L^{*}=0$,  $L^{\prime}\subset[-N,-1]$, and $L^{\prime\prime}\subset\sup L^{*}+[1,N]$, with the convention that $L^{\prime\prime}=\emptyset$ if $L^{*}$ is infinite (if $M$ is half-factorial, then $\Delta(M)=\emptyset$, $d=0$, $L^{\prime}=L^{\prime\prime}=\emptyset$, $L^{*}=\{0\}$, and $\mathcal{U}_{k}(M)=y_{k}+\{0\}=\{k\}$ for all $k\geq 1$).
We say that $M$ satisfies the Main Arithmetic Finiteness Properties if
  • (a)
    $M$ is locally tame.
  • (b)
    The set of distances is finite.
  • (c)
    The Structure Theorem for Unions holds.
  • (d)
    The Structure Theorem for Length Sets holds (it says that all length sets are almost arithmetic multiprogressions with global bounds for all parameters; for the technical definition see [64]).
Consider two simple examples of Mori domains, Cohen-Kaplansky domains and one-dimensional local Mori domains. Their arithmetic is extremely simple and it stimulates research to which extent the arithmetic of more general domains fulfills some resp. all of their arithmetic finiteness properties. Let $D$ be a domain and let $r$ be an ideal system on $D$ such that $\mathcal{I}_{r}(D)\subseteq\mathcal{I}(D)$. Then $D$ is called a Cohen-Kaplansky domain if it satisfies one of the following equivalent conditions ([71, Proposition 4.5])
  1. (a)
    $D$ is atomic and has only finitely many atoms up to associates.
  2. (b)
    The monoid of principal ideals is finitely generated.
  3. (c)
    The monoid $\mathcal{I}_{r}^{*}(D)$ of $r$-invertible $r$-ideals is finitely generated.
  4. (d)
    The semigroup $\mathcal{I}_{r}(D)$ of all nonzero $r$-ideals is finitely generated.
This class of domains was introduced by Anderson and Mott [4] (see also [36, 31]). Every Cohen-Kaplansky domain is one-dimensional, semilocal, Noetherian with nonzero conductor.
Theorem 11.3.
  1. (1)
    If $D$ is a one-dimensional local Mori domain, then $D$ satisfies the Main Arithmetic Finiteness Properties. Moreover, $M$ has finite elasticity if and only if the conductor is nonzero and the natural map $\operatorname{spec}(\widehat{D})\to\operatorname{spec}(D)$ is injective.
  2. (2)
    Let $M$ be a monoid such that $M_{{\text{\rm red}}}$ is finitely generated (e.g., the monoid of nonzero elements of a Cohen-Kaplansky domain). Then $M$ has finite elasticity and satisfies the Main Arithmetic Finiteness Properties.
Proof.
(1) The statement on the elasticity follows from [2, Theorem 2.12] and from [64, 2.9.2+2.10.7]. The other statements follow from [61, Theorem 4.1] and from [72, Theorem 3.9].
(2) This follows from [64, 3.1.4+4.4.11] and from [138]. ∎
12. Some Long-Term Arithmetic Problems
12.1. Half-Factoriality
In 1960 Carlitz proved that a ring of integers in an algebraic number field is half-factorial if and only if its class group has at most two elements. Ever since then half-factoriality is a key property studied in factorization theory and an abundance of results has been achieved (we refer to the surveys [29, 34] for some early work). We can pick only one recent highlight and one of our favorite problems.
After Carlitz’s result, the question which (non-principal) orders in number fields are half-factorial received a lot of attention. Many special cases were handled (including orders in quadratic number fields and seminormal orders), but only recently Rago [124] characterized half-factoriality for arbitrary orders in number fields.
In 1966 Claborn proved that every abelian group is isomorphic to the class group of a Dedekind domain. His result stimulated much further research. There is a lot of work on the realization of various types of abelian groups as class groups of Dedekind domains with wanted additional properties. We refer to some recent results and the literature cited there ([25, 117, 119]). The following problem was first posted by Steffan [137] in 1986. It has an affirmative answer for all Warfield groups [60] but is open in general.
PROBLEM 15.
Is every abelian group isomorphic to the class group of a half-factorial Dedekind domain?
12.2. Finiteness of Arithmetic Invariants
A key direction of research is to characterize (in a distinguished class of monoids) the finiteness of one resp. of all the arithmetic properties as they hold true in Theorem 11.3. We discuss a sample of results and then we give a more precise formulation of some problems. We start with a result by Kainrath [99].
Theorem 12.1.
Let $D$ be a domain that is finitely generated as a $\mathbb{Z}$-algebra. Then the elasticity $\rho(D)$ is finite if and only if $\mathcal{C}_{v}(\overline{D})$ and $\overline{D}/D$ are both finite and the natural map $\operatorname{spec}(\overline{D})\to\operatorname{spec}(D)$ is injective.
We continue with finiteness results for Krull monoids which highlight the significance of their $v$-class group for the arithmetic.
Theorem 12.2.
Let $M$ be a Krull monoid and let $G_{0}\subseteq\mathcal{C}_{v}(M)$ denote the set of classes containing prime divisors.
  1. (1)
    If $G_{0}$ is finite, then the Main Arithmetic Finiteness Properties hold and $\rho(M)<\infty$.
  2. (2)
    Suppose that $\mathcal{C}_{v}(M)$ is infinite cyclic, say $\mathcal{C}_{v}(M)=\mathbb{Z}$. Then $\rho(M)<\infty$ if and only if $\Delta(M)$ is finite if and only if all $\mathcal{U}_{k}(M)$’s are finite and the Structure Theorem for Unions holds if and only if either $G_{0}\cap\mathbb{N}$ or $G_{0}\cap(-\mathbb{N})$ is finite.
Proof.
The results of (1) can be found in [64] and (2) follows from [62, Theorem 4.2]. ∎
More is known than what is given by the above results. Among others, the assumption that $G_{0}$ is finite can be generalized to the assumption that the Davenport constant of $G_{0}$ is finite. If $\mathcal{C}_{v}(M)$ is finitely generated, then there is a characterization (in geometric terms of $G_{0}$) when $\rho(M)$ is finite ([82]). In particular, we do know that, in the above setting, length sets are completely controlled by the set $G_{0}$ (see the next subsection), but we are far away from a characterization of when one of the Main Arithmetic Finiteness Properties (a) - (d) hold true in general.
PROBLEM 16.
Let $M$ be a Krull monoid and let $G_{0}\subseteq\mathcal{C}_{v}(M)$ denote the set of classes containing prime divisors. Characterize (in terms of $G_{0}$) when one resp. all of the Main Arithmetic Finiteness Properties hold.
In the local case Theorem 11.3 provides a satisfactory answer for one-dimensional domains, but only little is known for higher-dimensional local domains. Apart from the local case and apart from Krull monoids, C-monoids are the most studied class of Mori monoids with respect to the arithmetic, and they include a class of higher-dimensional Mori domains (Theorem 12.4.2). We recall their definition.
Let $F$ be a factorial monoid and let $M\subseteq F$ be a submonoid. Two elements $y,y^{\prime}\in F$ are called $M$-equivalent (we write $y\sim y^{\prime}$) if $y^{-1}M\cap F={y^{\prime}}^{-1}M\cap F$ or, in other words,
$\text{if for all}\ x\in F,\ \text{we have}\ xy\in M\quad\text{if and only if}\quad xy^{\prime}\in M\,.$
$M$-equivalence defines a congruence relation on $F$ and for $y\in F$, let $[y]_{M}^{F}=[y]$ denote the congruence class of $y$. Then
$\mathcal{C}(M,F)=\{[y]\mid y\in F\}\quad\text{and}\quad\mathcal{C}^{*}(M,F)=\{[y]\mid y\in(F\setminus F^{\times})\cup\{1\}\}$
are commutative semigroups with identity element $[1]$. $\mathcal{C}(M,F)$ is the class semigroup of $M$ in $F$ and the subsemigroup $\mathcal{C}^{*}(M,F)\subseteq\mathcal{C}(M,F)$ is the reduced class semigroup of $M$ in $F$.
A monoid $M$ is a C-monoid if it is a submonoid of some factorial monoid $F$ such that $M^{\times}=M\cap F^{\times}$ and the reduced class semigroup is finite, and a domain $D$ is a C-domain if $D^{\bullet}$ is a C-monoid. The next theorem gathers the main algebraic and arithmetic properties of C-monoids (see [64]).
Theorem 12.3.
Let $M$ be a C-monoid.
  1. (1)
    $M$ is Mori, $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$, and $\mathcal{C}_{v}(\widehat{M})$ is finite.
  2. (2)
    $M$ satisfies the Main Arithmetic Finiteness Properties.
This result gives rise to the following two problems.
PROBLEM 17.
Let $M$ be a C-monoid. By Theorem 12.3.(2), we know that arithmetic invariants are finite but the present methods just provide abstract finiteness results. Thus the goal is to derive the precise values (or at least good upper bounds) for arithmetic invariants in terms of suitable algebraic invariants (say in terms of the class semigroup). Much is known is the special case when $M$ is Krull (see the next subsection). But in the general case of C-monoids only first steps are done so far (e.g., [68]).
PROBLEM 18.
Let $M$ be a Mori monoid with $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$ such that $\mathcal{C}_{v}(\widehat{M})$ is finite. Characterize when $M$ is a C-monoid. The cases when $M$ is seminormal or strong Mori are of special interest ([74]).
There is a simple characterization of C-monoids within the class of finitely generated monoids and a partial answer for Mori domains. But the general case is open.
Theorem 12.4.
  1. (1)
    Let $M$ be a monoid such that $M_{{\text{\rm red}}}$ is finitely generated. Then $M$ is Mori with $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$, and $M$ is a C-monoid if and only if $\mathcal{C}_{v}(\widehat{M})$ is finite.
  2. (2)
    Let $D$ be Mori domain with $(D\negthinspace:\negthinspace\widehat{D})\neq\{0\}$ and with $\mathcal{C}_{v}(\widehat{D})$ finite. If $\widehat{D}/(D\negthinspace:\negthinspace\widehat{D})$ is finite, then $D$ is a C-domain. The converse holds if $D$ is semilocal, Noetherian but not local.
Proof.
For Part (1), we refer to [66, Proposition 4.8]. The first statement in Part (2) follows from [64] and the converse statement is proved by Reinhart in [127, Corollary 4.5]. ∎
Every generalized Cohen-Kaplansky domain (this is an atomic domain that has only finitely many atoms, up to associates, that are not prime) is a C-domain ([71, Proposition 4.11]). Part 2 of Theorem 12.4 shows that, in particular, orders in algebraic number fields are C-domains. A classical result of Furtwängler characterizes the ideals that can be realized as conductor ideals of an order ([87, Theorem 2.12.10]). This was generalized by Reinhart to a more general setting of Noetherian domains ([128]), and his result gives rise to the following problem.
PROBLEM 19.
Let $H$ be a Krull monoid with finite $v$-class group. Characterize the ideals $\mathfrak{f}\subseteq H$ for which there is a C-monoid $M\subseteq H$ with $\widehat{M}=H$ and with conductor $\mathfrak{f}=(M\negthinspace:\negthinspace\widehat{M})$.
C-domains have been generalized to weakly C-domains which cover a larger class of Mori domains, as given in Theorem 12.4.2. They also satisfy most of the Main Arithmetic Finiteness Properties, but it is open whether the Structure Theorem for Length Sets holds true for them.
PROBLEM 20.
Let $M$ be a weakly C-monoid in the sense of [100]. Prove or disprove that the Structure Theorem for Length Sets holds (for a first step see [67]).
Based on a result of Nagata, Auslander and Buchsbaum proved that every regular local Noetherian domain is factorial. The following problem opens the door for a wide open area where only first steps have been done so far (e.g., [143]).
PROBLEM 21.
Study the interplay of homological properties (as discussed in the previous sections) and arithmetic finiteness properties of a domain.
12.3. The Transfer Krull Property
A monoid homomorphism $\theta\colon M\to B$, between atomic monoids $M$ and $B$, is called a transfer homomorphism if the following two properties hold.
  1. (T 1)
    $B=\theta(M)B^{\times}$ and $\theta^{-1}(B^{\times})=M^{\times}$.
  2. (T 2)
    If $u\in M$,  $b,\,c\in B$ and $\theta(u)=bc$, then there exist  $v,\,w\in M$ such that $u=vw$,   $\theta(v)\in bB^{\times}$ and $\theta(w)\in cB^{\times}$.
It is easy to check that $\mathsf{L}_{M}(a)=\mathsf{L}_{B}\big{(}\theta(a)\big{)}$ for all $a\in M$ and $\mathcal{L}(M)=\mathcal{L}(B)$. In particular, we have $\rho(M)=\rho(B)$ and $\Delta(M)=\Delta(B)$.
Let $M$ be a Krull monoid and let $G_{0}\subseteq\mathcal{C}_{v}(M)$ denote the set of classes containing prime divisors. Then there is a transfer homomorphism $\theta\colon M\to\mathcal{B}(G_{0})$, where $\mathcal{B}(G_{0})$ is the monoid of zero-sum sequences with terms from $G_{0}$. This allows to study arithmetic invariants of $M$ in the combinatorial monoid $\mathcal{B}(G_{0})$, which is studied with methods from additive combinatorics. The interplay is most strong if every class contains a prime divisor (so, if $G_{0}=\mathcal{C}_{v}(M)$) and $\mathcal{C}_{v}(M)$ is finite, say $\mathcal{C}_{v}(M)\cong C_{n_{1}}\oplus\ldots\oplus C_{n_{r}}$ with $1<n_{1}\mid\ldots\mid n_{r}$. We refer to [132, 63] which demonstrates that arithmetic invariants of $M$ can be studied in terms of the group invariants $(n_{1},\ldots,n_{r})$.
Along the way, it was observed that there are monoids and domains, that are not Krull, but nevertheless they allow transfer homomorphisms to monoids of zero-sum sequences. Inspired by the success of the powerful methods from additive combinatorics, intense investigations have been started to figure out under which conditions we do have transfer homomorphisms to monoids of zero-sum sequences. Deep results revealed that this may happen even for non-commutative rings ([136]). Here we restrict our discussion to monoids (so, to commutative cancellative semigroups) and just mention two results. A one-dimensional local Mori domain $D$ is transfer Krull if and only if $D$ is half-factorial (whence in the trivial case only; see [73, Theorem 5.5]). Recent characterizations for orders in Dedekind domains are given in [125]. For a survey and more, see [15, 75, 14].
PROBLEM 22.
Take a distinguished class of Mori monoids and characterize (in terms of algebraic invariants) when they are transfer Krull.
Module theory offers a rich source of not necessarily cancellative semigroups of modules that are transfer Krull (for such an example, see [12]). We formulate the following problem.
PROBLEM 23.
Let $R$ be a not necessarily commutative ring and let $\mathcal{C}$ denote a class of left $R$-modules, that is closed under direct sums, direct summands, and isomorphisms. Let $\mathcal{V}(\mathcal{C})$ be the semigroup of isomorphism classes of modules from $\mathcal{C}$, with operation induced by the direct sum. If every module from $\mathcal{C}$ has a semilocal endomorphism ring, then $\mathcal{V}(\mathcal{C})$ is Krull (this was proved by Facchini; for examples, see [44]). For a given class $\mathcal{C}$, for which $\mathcal{V}(\mathcal{C})$ is Krull, study classes $\mathcal{C}^{\prime}$ of modules having a transfer homomorphism $\theta\colon\mathcal{V}(\mathcal{C}^{\prime})\to\mathcal{V}(\mathcal{C})$. For a simple (artificial) example, of how this strategy could work, see [75, Example 5.4].
12.4. What can a family of length sets look like?
In contrast to Theorem 12.2 there is the following result by Kainrath (see [64, Theorem 7.4.1]).
Theorem 12.5.
Let $M$ be a Krull monoid and let $G_{0}\subseteq\mathcal{C}_{v}(M)$ denote the set of classes containing prime divisors. If $G_{0}$ contains an infinite abelian group, then every finite, nonempty subset of $\mathbb{N}_{\geq 2}$ occurs as a length set of $M$.
The same result is true for large classes of integer-valued polynomials over Krull domains and of weakly Krull Mori domains (for a sample, see [24, 46, 47]). Surprisingly in almost all settings studied so far, monoids either satisfy the Structure Theorem for Length Sets or every finite, nonempty subset of $\mathbb{N}_{\geq 2}$ occurs as a length set. This dichotomy gave rise to the following problem. Let $\mathcal{P}_{\text{\rm fin}}(\mathbb{N}_{0})$ denote the system of all finite, nonempty subsets of $\mathbb{N}_{0}$.
PROBLEM 24.
Which families $\mathcal{L}\subseteq\mathcal{P}_{\text{\rm fin}}(\mathbb{N}_{0})$ can be realized as the system of length sets of a Mori monoid?
Let $\mathcal{L}\subseteq\mathcal{P}_{\text{\rm fin}}(\mathbb{N}_{0})$ be a family of finite, nonempty subsets. If $\mathcal{L}$ is the system of length sets of a Mori monoid, which is not a group, then the following properties hold.
  • (i)
    $\{0\},\{1\}\in\mathcal{L}$ and all other sets of $\mathcal{L}$ lie in $\mathbb{N}_{\geq 2}$.
  • (ii)
    For every $k\in\mathbb{N}_{0}$ there is $L\in\mathcal{L}$ with $k\in L$.
  • (iii)
    If $L_{1},L_{2}\in\mathcal{L}$, then there is $L\in\mathcal{L}$ with $L_{1}+L_{2}\subseteq L$.
If $\mathcal{L}$ is a family satisfying the stronger condition that for each two $L_{1},L_{2}\in\mathcal{L}$ the sumset $L_{1}+L_{2}\in\mathcal{L}$, then there is a Dedekind domain $D$ such that $\mathcal{L}(D)=\mathcal{L}$ ([76]). But the general case is open.
The system $\mathcal{L}(M)$ gathers all length sets of a given monoid $M$ and Theorems 12.2 and 12.5 describe all the phenomena that might occur. But, in addition, one might ask after the typical structure of a length set when we pick a random element. Let $M$ be an analytic monoid. This is a Krull monoid with finite $v$-class group and with a norm function that allows to establish a theory of L-functions ([98]). Then almost all (in the sense of Dirichlet density) length sets are intervals ([64, Theorem 9.4.11]). Other concepts of density include the Zariski density (say for $\mathbb{Q}$-algebras). There is much heuristic evidence (in a variety of settings) that most (in what sense ever) length sets are much simpler compared with the possible structure of all length sets popping up in $\mathcal{L}(M)$. Thus, we present the following problem.
PROBLEM 25.
Consider a distinguished class of Mori monoids which allows us to establish a natural concept of density. Describe the structure of length sets of subsets of elements having high density.
We end with a short discussion on semigroups of ideals that are not necessarily cancellative. The algebraic and arithmetic structure of ideal semigroups has found wide attention in the literature (for a sample of recent contributions, see [96, 3, 48, 71, 18, 97, 69, 58, 108]). In general, these semigroups need not be cancellative (indeed, the semigroup of all nonzero ideals of a domain is cancellative if and only if the domain is almost Dedekind [84, Theorem 23.2]). However, semigroups of ideals are unit-cancellative in a variety of significant cases. To discuss this situation, let $M$ be a commutative semigroup with identity. We say that $M$ is unit-cancellative if $a,u\in M$ and $au=a$ implies that $u\in H^{\times}$. Factorization theory in unit-cancellative monoids can be developed along the same lines as in the cancellative setting. In particular, atoms and length sets are defined as it is done for monoids in Section 2.
Let $D$ be a domain and let $r$ be an ideal system on $D$. Suppose that $D$ satisfies the $r$-Krull Intersection Theorem, which means that
$\bigcap_{n\in\mathbb{N}_{0}}(I^{n})_{r}=\{0\}\quad\text{for all $I\in\mathcal{I}_{r}(D)\setminus\{D\}$.}$
Now let $I,J\in\mathcal{I}_{r}(D)$ with $(IJ)_{r}=I$. This implies that $(IJ^{n})_{r}=I$ for all $n\in\mathbb{N}$, whence $\{0\}\neq I\subseteq\cap_{n\in\mathbb{N}}(J^{n})_{r}$. Since $D$ satisfies the $r$-Krull Intersection Theorem, we infer that $J=D$, whence $\mathcal{I}_{r}(D)$ is unit-cancellative.
PROBLEM 26.
Let $D$ be a domain and let $r$ be an ideal system on $D$ such that $\mathcal{I}_{r}(D)$ is unit-cancellative. Study the system of length sets of $\mathcal{I}_{r}(D)$.
13. Appendix: Non-Commutative Rings
Multiplicative ideal theory and factorization theory spread from commutative ring theory to noncommutative ring theory in the first half of the 20th century. Although this is beyond the scope of our present article, we would be remiss not even to mention this development.
A starting point of ideal theory in the non-commutative setting was the pioneering work of Asano [5], followed by work of Asano, Murato, and Harada [6, 89], which led to a theory of Dedekind-like rings (including Asano prime rings and Dedekind prime rings, see [112]). From the 1970s on, non-commutative Krull rings have been introduced ([113, 23, 95]), and it was Wauters who introduced non-commutative Krull monoids in order to study non-commutative semigroup algebras ([142, 94, 59]). We refer to some recent articles, with a focus on surveys [129, 1, 38, 131, 130].
The early development of factorization theory in non-commutative rings is discussed in a survey by Smertnig [135]. We briefly address two current research directions.
As it is the case in commutative ring theory, a topic of research are the questions which atomic prime rings have the finite factorization property (meaning that each element has only finitely many essentially different factorizations) resp. the bounded factorization property (meaning that all length sets are finite and nonempty; see [17, 16]). Whereas it is quite easy to verify that a commutative Noetherian domain has bounded factorizations (see Lemma 11.2), it is still an open problem whether every Noetherian prime ring has bounded factorizations.
A further key problem asks which Noetherian prime rings allow a transfer homomorphism to a commutative (cancellative) monoid or even to a commutative Krull monoid. In his seminal paper [134], Smertnig characterized maximal orders in central simple algebras over algebraic number fields, that allow a transfer homomorphism to a commutative Krull monoid (the proof is based on theory of one-sided divisorial ideals in such orders). This result initiated much further research (e.g., [10, 11, 136, 13, 9]). Note that the existence of a transfer homomorphism to a commutative Krull monoid immediately implies the existence of a transfer homomorphism to a monoid of zero-sum sequences over a subset of an abelian group, which in turn opens the door for methods from additive combinatorics (see the Subsection on the Transfer Krull Property).
Acknowledgement. We would like to thank our colleagues Valentin Havlovec, Giulio Peruginelli, Andreas Reinhart, and Daniel Smertnig for their feedback and all their helpful comments.

### v3

@@
-On Long-Term Problems in Multiplicative Ideal Theory and Factorization Theory
+arXiv:2502.21020v3[math.AC] 28 Feb 2026
 Abstract.
@@
 Let $\mathbb{N}$ denote the set of positive integers and set $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$.
-For integers $a,b\in\mathbb{Z}$, let $[a,b]=\{x\in\mathbb{Z}\colon a\leq x\leq b\}$ denote the discrete interval between $a$ and $b$.
+For integers $a,b\in\mathbb{Z}$, let $a,b]=\{x\in\mathbb{Z}\colon a\leq x\leq b\}$ denote the discrete interval between $a$ and $b$.
 For subsets $A,B$ of an additive abelian group $G$, let $A+B=\{a+b\colon a\in A,b\in B\}\subseteq G$ denote their sumset.
@@
 We denote by $\mathcal{I}_{r}(M)$ the set of all nonempty $r$-ideals, and we define $r$-multiplication by setting $I\cdot_{r}J=(IJ)_{r}$ for all $I,J\in\mathcal{I}_{r}(M)$.
-Then $\mathcal{I}_{r}(M)$ together with $r$-multiplication is a reduced semigroup with identity element $M$.
+Then $\mathcal{I}_{r}(M)$ together with $r$-multiplication is a reduced semigroup with identity element $M$, and it is a multiplicative lattice.
+Conversely, by [39], for every multiplicative lattice there are a commutative semigroup $H$ and a weak ideal system $r$ on $H$ such that $L$ is isomorphic to the lattice of $r$-ideals of $H$.
 A subset $X\subseteq\mathsf{q}(M)$ is called ($M$-)fractional if $cX\subseteq M$ for some $c\in M$. We define
@@
 Wang (see Section 7).
-All rings are commutative and have an identity element.
-For a commutative ring $D$, we denote by $D^{\times}$ its group of units and by $D^{\bullet}$ its monoid of regular elements.
-Let $D$ be a domain with quotient field $K$.
-We denote by $\mathcal{A}(D)$ the set of atoms, by $\mathsf{q}(D)$ the quotient field of $D$, by $\overline{D}$ the integral closure of $D$, and by $\widehat{D}=\widehat{D^{\bullet}}\cup\{0\}$ the complete integral closure of $D$.
+[…]
 An ideal system on $D$ is a map $r\colon\mathcal{P}(D)\to\mathcal{P}(D)$ such that the following conditions are satisfied for all subsets $X,Y\subseteq D$ and all $c\in D$:
@@
 For a polynomial $f\in V[X]$, let $c(f)$ be the ideal of $V$ generated by the coefficients of $f$.
-Let $N=\{f\in V[X]\mid c(f)=V\}$ and finally, let $V(X)=N^{-1}V[X]$.
-$V(X)$ is known as the Nagata extension of $V$.
+Let $N=\{f\in V[X]\mid c(f)=V\}$ and finally, let $N^{-1}V[X].
+V(X)V(X)$ is known as the Nagata extension of $V$.
 $V(X)$ is a valuation domain which is almost identical to $V$.
@@
 4. Classification of Integral Domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$
-We survey recent developments in the classification of domains that lie strictly between the polynomial ring $\mathbb{Z}[X]$ and its field of fractions $\mathbb{Q}[X]$. We focus on two major studies:
+We survey recent developments in the classification of domains that lie strictly between the polynomial ring $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$. We focus on two major studies:
   • [111]
@@
   • $\mathbb{C}_{p}$ denotes the completion of the algebraic closure of the field $\mathbb{Q}_{p}$.
-  • ¯$\overline{\mathbb{Z}}_{p}$ denotes the absolute integral closure of $\mathbb{Z}_{p}$.
-  • ^¯$\overline{\widehat{\mathbb{Z}}}=\prod_{p\in\mathbb{P}}\overline{\mathbb{Z}}_{p}$.
+  • $\overline{\mathbb{Z}}_{p}$ denotes the absolute integral closure of $\mathbb{Z}_{p}$.
+  • $\overline{\widehat{\mathbb{Z}}}=\prod_{p\in\mathbb{P}}\overline{\mathbb{Z}}_{p}$.
 A generalized ring of integer-valued polynomials is defined, for a subset $\underline{E}\subset\overline{\widehat{\mathbb{Z}}}$, as
@@
 Let $\underline{E}=\prod_{p}E_{p}$ be a subset of $\overline{\widehat{\mathbb{Z}}}$.
-Following [117, Definition 2.11], we say that $\underline{E}$ is polynomially factorizable if, for every $g\in\mathbb{Z}[X]$ and every $\alpha=(\alpha_{p})\in\underline{E}$, there exist $n,d\in\mathbb{Z}$ with $n,d\geq 1$ such that $g(\alpha)^{n}/d$ is a unit of $\overline{\widehat{\mathbb{Z}}}$.
+Following [118, Definition 2.11], we say that $\underline{E}$ is polynomially factorizable if, for every $g\in\mathbb{Z}[X]$ and every $\alpha=(\alpha_{p})\in\underline{E}$, there exist $n,d\in\mathbb{Z}$ with $n,d\geq 1$ such that $g(\alpha)^{n}/d$ is a unit of $\overline{\widehat{\mathbb{Z}}}$.
 In [118], the author classifies Dedekind domains with finite residue fields of prime characteristic that lie between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$.
@@
 Define $\mathcal{O}=\prod_{p\in\mathbb{P}}\mathbb{O}_{p}\subset\prod_{p\in\mathbb{P}}\mathbb{C}_{p}$.
-Theorem 4.2 ([118, Theorem 3.4]).
+Theorem 4.2 ([119, Theorem 3.4]).
 Let $R$ be a Dedekind domain such that $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$.
@@
 This result completes the classification of Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ and provides a powerful characterization of Dedekind domains in this setting, linking local valuation data and integer-valued polynomial constraints to global structural properties such as the class group.
-The result is obtained by characterizing discrete valuation rings (DVRs) $W$ of the field $\mathbb{Q}(X)$ that are residually algebraic over $\mathbb{Z}_{(p)}$ for some prime $p$ (see [118, Corollary 2.28]).
+The result is obtained by characterizing discrete valuation rings (DVRs) $W$ of the field $\mathbb{Q}(X)$ that are residually algebraic over $\mathbb{Z}_{(p)}$ for some prime $p$ (see [119, Corollary 2.28]).
 When the extension of residue fields has infinite degree, such a domain is given by
@@
 In particular, unifying generalized integer-valued polynomial viewpoints with valuation-theoretic frameworks remains an active area of research, as do possible applications to:
-  • Krull or ((((strong)))) Mori domains: Extending classification results to settings with additional constraints (e.g., finite residue fields, specific $v$-class groups, etc.).
+  • Krull or ((strong)) Mori domains: Extending classification results to settings with additional constraints (e.g., finite residue fields, specific $v$-class groups, etc.).
   • Factorial domains: Investigating when generalized integer-valued polynomial rings are factorial.
@@
   • The actual construction is built using the ab star operations. It is cumbersome. One could wish for a simpler construction.
-  • One could wonder whether the base ring $D$ is essential. Is it possible to build a Kronecker function domain using just $K$ without $D$?
+  • One could wonder whether the base ring $D$ is essential. Is it possible to build a Kronecker function domain using just $K$ - without $D$?
 In 2003 Halter-Koch [87] gave what he called an axiomatic treatment of what he called Kronecker domains which gives us the construction we were looking for above.
@@
 Halter-Koch’s notion of Kronecker domain has been known since 2003 and very little has been done in that time to study them.
-Actually, Halter-Koch himself extended what is known by himself by demonstrating that they can be built with a variant of the conventional construction ([88, Theorem 6.5.5]).
+Actually, Halter-Koch himself extended what is known by himself by demonstrating that they can be built with a variant of the conventional construction ([90, Theorem 6.5.5]).
 Many questions could be posed.
@@
 On the other hand, $M$ is called GV-torsion if, for every nonzero element $x\in M$, there exists $J\in\operatorname{GV}(R)$ such that $Jx=0$.
-It is evident that the pair (GV-torsion modules, GV-torsion-free modules) defines a hereditary torsion theory, first introduced in [102, p.
+It is evident that the pair (GV-torsion modules, GV-torsion-free modules) defines a hereditary torsion theory, first introduced in [104, p.
 1618].
@@
 Using $w$-operations, strong Mori domains were introduced in [141].
-It was previously claimed ([57, Corollary IX.5.5]) that for a domain $D$ with quotient field $K$, the injective module $E(K/D)$ is $\Sigma$-injective if and only if $D$ is a Mori domain.
-However, this statement is incorrect: one must replace ‘Mori’ with ‘strong Mori’ (see [56, Theorem 9]).
+It was previously claimed ([59, Corollary IX.5.5]) that for a domain $D$ with quotient field $K$, the injective module $E(K/D)$ is $\Sigma$-injective if and only if $D$ is a Mori domain.
+However, this statement is incorrect: one must replace ‘Mori’ with ‘strong Mori’ (see [58, Theorem 9]).
 In recent years, many classical theorems about Noetherian rings have been generalized to strong Mori domains.
@@
     Every nonzero non-unit in $D$ is contained in only finitely many height-one prime ideals.
-Analogous to how Dedekind domains can be characterized by the invertibility of nonzero ideals, Krull domains can be characterized as follows ([101, Theorem 3.6], [139, Theorem 7.9.3]):
+Analogous to how Dedekind domains can be characterized by the invertibility of nonzero ideals, Krull domains can be characterized as follows ([103, Theorem 3.6], [140, Theorem 7.9.3]):
 […]
@@
   2. (b)
-    The $w$-factor ring $(D/I)_{w}$ is a zero-dimensional semilocal principal ideal ring for every proper $w$-ideal $I$ of $D$, [26, Theorem 4.5].
+    The $w$-factor ring $(D/I)_{w}$ is a zero-dimensional semilocal principal ideal ring for every proper $w$-ideal II of $D$, [26, Theorem 4.5].
   3. (c)
-    $(D/I)_{w}$ is a quasi-Frobenius ring [148, Theorem 9].
-[…]
+    $(D/I)_{w}$ is a quasi-Frobenius ring [149, Theorem 9].
+These results highlight a strong parallel between Dedekind and Krull domains when analyzed through $w$-operations and $w$-ideal structures.
+This raises the question of whether PvvMDs can be characterized by their $w$-factor rings, where an integral domain $D$ is called a Prüfer $v$-multiplication domain (PvvMD for short) if every nonzero finitely generated ideal of $D$ is $t$-invertible (or $w$-invertible).
 PROBLEM 8.
-[…]
+Investigate whether the PvvMDs $D$ are precisely those domains for which their $w$-factor rings $(D/I)_{w}$ are FC-rings for every proper $w$-ideal II of $D$, where a commutative ring $R$ is an FC-ring if it is coherent and self-FP-injective.
 8. Homological versus Elementary Methods
@@
     $\cdots\;\longrightarrow\;F_{1}\;\longrightarrow\;F_{0}\;\longrightarrow\;F^{0}\;\longrightarrow\;F^{1}\;\longrightarrow\;\cdots$
-    such that $M\cong\ker(F^{0}\to F^{1})$ and exactness is preserved under the functor $-\otimes_{R}E$ for every injective $R$-module $E$.
+    such that $M\cong\ker(F^{0}\to F^{1})$ and exactness is preserved under the functor $\otimes_{R}E$ for every injective $R$-module $E$.
   4. (4)
@@
 Yet there remain instances in which Gorenstein analogies are only partially true or require stricter hypotheses.
-For example, in the classical setting, a domain $D$ is Prüfer if and only if it is locally a valuation domain ([139, Corollary 3.7.22]).
-In the Gorenstein setting, while every G-Prüfer domain is locally G-Prüfer [145, Corollary 11], the converse fails in general [147, Example 4.1].
+For example, in the classical setting, a domain $D$ is Prüfer if and only if it is locally a valuation domain ([140, Corollary 3.7.22]).
+In the Gorenstein setting, while every G-Prüfer domain is locally G-Prüfer [146, Corollary 11], the converse fails in general [148, Example 4.1].
 Hence, the direct classical analogy does not hold perfectly in Gorenstein contexts.
-Classically, $D$ is Prüfer if and only if every overring of $D$ is flat [107, Theorem 6.10].
-Analogously, it has been shown that if $R$ is G-Prüfer, then every overring of $R$ is Gorenstein flat [145, Theorem 6].
+Classically, $D$ is Prüfer if and only if every overring of $D$ is flat [109, Theorem 6.10].
+Analogously, it has been shown that if $R$ is G-Prüfer, then every overring of $R$ is Gorenstein flat [146, Theorem 6].
 However, it remains open whether the converse also holds.
@@
 This is a natural Gorenstein-homological parallel to a well-known Prüfer criterion, but it awaits resolution.
-A key feature of Dedekind domains $D$ is that every nonzero ideal is invertible, which is equivalent to demanding that $D/I$ be quasi-Frobenius for each nonzero ideal $I$ ([148, Theorem 13]).
+A key feature of Dedekind domains $D$ is that every nonzero ideal is invertible, which is equivalent to demanding that $D/I$ be quasi-Frobenius for each nonzero ideal $I$ ([149, Theorem 13]).
 Similarly, in the Gorenstein Dedekind setting, one has:
-  • An integral domain $D$ is G-Dedekind if and only if, for every nonzero element $a\in D$, the factor ring $D/(a)$ is quasi-Frobenius ([139, Theorem 11.7.7]).
+  • An integral domain $D$ is G-Dedekind if and only if, for every nonzero element $a\in D$, the factor ring $D/(a)$ is quasi-Frobenius ([140, Theorem 11.7.7]).
 This implies that every nonzero ideal of $D$ is a $v$-ideal.
 In fact, $D$ is a G-Dedekind domain if and only if it is a Noetherian divisorial domain.
-  • In general, a G-Dedekind domain is one-dimensional and Noetherian ([139, Theorem 11.7.7 and Corollary 11.7.8]), yet can exhibit different behaviors regarding integral closure, reminiscent of how Gorenstein rings relax some classical regularity or integrality conditions.
-Recent connections show that factor rings of the form $(D/xD)_{w}$ being quasi-Frobenius for all nonzero $x$ in $D$ characterizes G-Krull domains [146, Theorem 2.7].
+  • In general, a G-Dedekind domain is one-dimensional and Noetherian ([140, Theorem 11.7.7 and Corollary 11.7.8]), yet can exhibit different behaviors regarding integral closure, reminiscent of how Gorenstein rings relax some classical regularity or integrality conditions.
+Recent connections show that factor rings of the form $(D/xD)_{w}$ being quasi-Frobenius for all nonzero $x$ in $D$ characterizes G-Krull domains [147, Theorem 2.7].
 Such Gorenstein-enabled statements suggest deep analogies between multiplicative ideal theory and Gorenstein homological algebra, but also underscore certain exceptions where classical and Gorenstein theories diverge.
@@
     and
-    =$\mathcal{B}\;=\;\bigl{\{}\,N\;\bigm{|}\;\operatorname{Ext}^{1}_{R}(A,N)=0\text{ for all }A\in\mathcal{A}\bigr{\}}.$
+    $\mathcal{B}\;=\;\bigl\{\,N\;\bigm|\;\operatorname{Ext}^{1}_{R}(A,N)=0\text{ for all }A\in\mathcal{A}\bigr\}.$
   2. (2)
@@
   2. (2)
-    Over a principal ideal domain $D$, the cotorsion pair $(\mathcal{A},\mathcal{B})=(\mathrm{Flat},\mathrm{Cotorsion})$ satisfies:
+    Over any ring $D$, the cotorsion pair $(\mathcal{A},\mathcal{B})=(\mathrm{Flat},\mathrm{Cotorsion})$ satisfies:
     $\mathcal{A}\;=\;\{\text{flat }D\text{-modules}\},\quad\mathcal{B}\;=\;\bigl\{\,N\;\bigm|\;\operatorname{Ext}^{1}_{R}(F,N)=0\text{ for all flat }F\bigr\}.$
@@
 By focusing on how $\operatorname{Ext}^{1}_{R}(-,-)$ vanishes, one can obtain powerful decomposition theorems for modules.
-A variant of cotorsion theory, called Tor-cotorsion theory, shifts attention from $\operatorname{Ext}$ to $\operatorname{Tor}$ and hence emphasizes torsion properties and tensor products.
+A variant of cotorsion theory, called Tor-torsion theory, shifts attention from $\operatorname{Ext}$ to $\operatorname{Tor}$ and hence emphasizes torsion properties and tensor products.
 Definition 10.3.
-An $R$-module $M$ is called Tor-cotorsion if
-$\operatorname{Tor}_{1}^{R}(F,M)\;=\;0\quad\text{for every flat $R$-module }F.$
+Let $\mathcal{A}$ and $\mathcal{B}$ be two classes of modules. The pair $(\mathcal{A},\mathcal{B})$ is called a Tor-torsion pair (or Tor-torsion theory) if
+$\mathcal{A}=\mathcal{B}^{\top}\quad\text{and}\quad\mathcal{B}=\mathcal{A}^{\top},$
+where
+$\mathcal{A}^{\top}=\left\{B\in\mathcal{M}\middle|\operatorname{Tor}_{1}^{R}(A,B)=0\text{ for every }A\in\mathcal{A}\right\}.$
 In classical cotorsion theory, one requires $\operatorname{Ext}^{1}_{R}(F,M)=0$ for every flat $R$-module $F$.
-In Tor-cotorsion theory, one instead demands that $\operatorname{Tor}_{1}^{R}(F,M)=0$ for every flat $R$-module $F$.
+In Tor-torsion theory, one instead demands that $\operatorname{Tor}_{1}^{R}(F,M)=0$ for every flat $R$-module $F$.
 Both conditions provide different perspectives on how flat modules (or certain torsion-free classes) might interact with a given module, illuminating complementary aspects of homological dimension and torsion phenomena.
@@
 Let $\mathscr{F}$ be the class of all flat $R$-modules.
-Then $(\mathscr{F},\mathcal{M})$ is a (classical) Tor-cotorsion pair in the category of $R$-modules.
-In a Tor-cotorsion context, we focus on $R$-modules $M$ with $\operatorname{Tor}_{1}^{R}(F,M)=0$ rather than on $\operatorname{Ext}^{1}_{R}(F,M)=0$.
-Cotorsion theory (and its Tor-cotorsion counterpart) can effectively describe structural aspects of rings that are central in multiplicative ideal theory. For instance:
+Then $(\mathscr{F},\mathcal{M})$ is a (classical) Tor-torsion pair in the category of $R$-modules.
+In a Tor-torsion context, we focus on $R$-modules $M$ with $\operatorname{Tor}_{1}^{R}(F,M)=0$ rather than on $\operatorname{Ext}^{1}_{R}(F,M)=0$.
+Cotorsion theory (and its Tor-torsion counterpart) can effectively describe structural aspects of rings that are central in multiplicative ideal theory. For instance:
   1. (1)
@@
   2. (2)
-    An integral domain $D$ is Krull if and only if the global weak $w$-projective dimension $w.w.gl.\dim(D)$ is at most $1$; see [141, Theorem 4.3].
+    An integral domain $D$ is Krull if and only if the global weak $w$-projective dimension $w.w.gl.\dim(D)$ is at most $1$; see [142, Theorem 4.3].
 These homological characterizations blend naturally with cotorsion-theoretic viewpoints: the condition $gl.\dim(R)\leq 1$ can be recast via orthogonality requirements on $R$-modules, while the $w$-projective setting relies on extension functors in the $w$-module theory sense.
@@
     $M$ is called a Lee cotorsion module if
-[…]
+    $\operatorname{Ext}^{1}_{D}(K/D,M)\;=\;0.$
 Theorem 10.6 ([140], Theorem 12.9.12).
@@
 This result interlinks classical ideals-based characterizations (e.g., localizing at height-one primes or checking invertibility of ideals) with module-theoretic vanishing conditions.
-[…]
+The interplay between cotorsion theory and $w$-theory highlights how star operations can refine homological statements, particularly for non-Noetherian rings or those with specialized divisorial properties.
+For example, focusing on $w$-flatness and $w$-cotorsion can unify certain results in multiplicative ideal theory with relative homological vanishing conditions.
+In this direction, the cases of PvvMDs and Krull domains have been studied in [122, 123].
 PROBLEM 14.
-Develop explicit cotorsion and Tor-cotorsion characterizations of rings and integral domains that arise in multiplicative ideal theory. In particular, investigate how homological invariants such as
-[…]
+Develop explicit cotorsion and Tor-torsion characterizations of rings and integral domains that arise in multiplicative ideal theory. In particular, investigate how homological invariants such as
+$\mathrm{Ext}^{1}_{R}(-,-),\quad\mathrm{Tor}_{1}^{R}(-,-),\quad\text{global dimensions, etc.}$
 can be used to classify such rings. Examples of interest include:
@@
     (Tor-)cotorsion approaches.
-[…]
+  • Other star-operation domains: Characterize specialized classes (e.g., Gorenstein domains, G-Prüfer, PvvMDs, etc.) via novel cotorsion or Tor-torsion properties.
 11. Factorization Theory in Mori Monoids
@@
 We start with a lemma highlighting the abundance of Mori and Krull monoids in commutative ring theory.
-For more on the ideal theory of Mori monoids we refer to [88, Chapter 5] and for more of this flavor in the setting of $v$-Marot and $t$-Marot rings see [27, 14].
+For more on the ideal theory of Mori monoids we refer to [90, Chapter 5] and for more of this flavor in the setting of $v$-Marot and $t$-Marot rings see [27, 14].
 For Mori monoids with a combinatorial flavor see [44].
 Moreover, we refer to a survey of Facchini [45] for the role of Krull monoids in module theory.
-skip
 Lemma 11.1.
@@
     Let $M$ be a Mori monoid.
-If $M$ is seminormal or if the conductor $(M\colon\widehat{M})$ is nonempty, then the complete integral closure $\widehat{M}$ is Krull.
+If $M$ is seminormal or if the conductor (M:M^)($\colon\widehat{M}$) is nonempty, then the complete integral closure $\widehat{M}$ is Krull.
 Furthermore, the complete integral closure of a strong Mori domain is Krull.
 Proof.
-For (1) and (2) we refer to [70] and for (3) to [69, Theorem 4.5]. The first claim in (4) is the Theorem of Mori-Nagata and the second claim follows from [65, Example 2.1].
-(5) For a seminormal $M$ the claim follows from [88, Theorem 5.2.8] and if the conductor is nonempty, then it follows from [64, Theorem 2.3.5].
-If $M$ is a strong Mori domain, then the claim follows from [139, Theorems 7.7.14 and 7.10.21].
+For (1) and (2) we refer to [72] and for (3) to [71, Theorem 4.5]. The first claim in (4) is the Theorem of Mori-Nagata and the second claim follows from [67, Example 2.1].
+(5) For a seminormal $M$ the claim follows from [90, Theorem 5.2.8] and if the conductor is nonempty, then it follows from [66, Theorem 2.3.5].
+If $M$ is a strong Mori domain, then the claim follows from [140, Theorems 7.7.14 and 7.10.21].
 ∎
@@
   1. (1)
-    All length sets are finite and nonempty (((([64, Theorem 2.2.9])))).
+    All length sets are finite and nonempty (([66, Theorem 2.2.9])).
   2. (2)
-    $\omega(a)<\infty$ for all $a\in M$ (((([65, Theorem 4.2])))).
+    $\omega(a)<\infty$ for all $a\in M$ (([67, Theorem 4.2])).
 For a finite, nonempty set $L=\{m_{1},\ldots,m_{k}\}\subset\mathbb{N}$ with $k\in\mathbb{N}$ and $m_{1}<\ldots<m_{k}$, we call
@@
 Thus, by definition,
-$M\ \text{is half-factorial}\quad\text{if and only if}\quad\Delta(M)=\emptyset\quad\text{if and only if}\quad\rho(M)=1\,,$
+[…]
 and, by a straightforward calculation, we get
@@
 Let $D$ be a domain and let $r$ be an ideal system on $D$ such that $\mathcal{I}_{r}(D)\subseteq\mathcal{I}(D)$.
-Then $D$ is called a Cohen-Kaplansky domain if it satisfies one of the following equivalent conditions ([71, Proposition 4.5])
+Then $D$ is called a Cohen-Kaplansky domain if it satisfies one of the following equivalent conditions ([73, Proposition 4.5])
   1. (a)
@@
 Proof.
-(1) The statement on the elasticity follows from [2, Theorem 2.12] and from [64, 2.9.2+2.10.7]. The other statements follow from [61, Theorem 4.1] and from [72, Theorem 3.9].
-(2) This follows from [64, 3.1.4+4.4.11] and from [138]. ∎
+(1) The statement on the elasticity follows from [2, Theorem 2.12] and from [66, 2.9.2+2.10.7]. The other statements follow from [63, Theorem 4.1] and from [74, Theorem 3.9].
+(2) This follows from [66, 3.1.4+4.4.11] and from [139]. ∎
 12. Some Long-Term Arithmetic Problems
@@
 In 1960 Carlitz proved that a ring of integers in an algebraic number field is half-factorial if and only if its class group has at most two elements.
-Ever since then half-factoriality is a key property studied in factorization theory and an abundance of results has been achieved (we refer to the surveys [29, 34] for some early work).
+Ever since then half-factoriality is a key property studied in factorization theory and an abundance of results has been achieved (we refer to the surveys [29, 34] for some early work and to [38] for a recent characterization of half-factoriality).
 We can pick only one recent highlight and one of our favorite problems.
@@
 Proof.
-The results of (1) can be found in [64] and (2) follows from [62, Theorem 4.2]. ∎
+The results of (1) can be found in [66] and (2) follows from [64, Theorem 4.2]. ∎
 More is known than what is given by the above results.
@@
 $\text{if for all}\ x\in F,\ \text{we have}\ xy\in M\quad\text{if and only if}\quad xy^{\prime}\in M\,.$
-$M$-equivalence defines a congruence relation on $F$ and for $y\in F$, let $[y]_{M}^{F}=[y]$ denote the congruence class of $y$. Then
+$M$-equivalence defines a congruence relation on $F$ and for $y\in F$, let $y]_{M}^{F}=[y]$ denote the congruence class of $y$. Then
 $\mathcal{C}(M,F)=\{[y]\mid y\in F\}\quad\text{and}\quad\mathcal{C}^{*}(M,F)=\{[y]\mid y\in(F\setminus F^{\times})\cup\{1\}\}$
@@
   1. (1)
-    $M$ is Mori, $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$, and $\mathcal{C}_{v}(\widehat{M})$ is finite.
+[…]
   2. (2)
@@
 PROBLEM 18.
-Let $M$ be a Mori monoid with $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$ such that $\mathcal{C}_{v}(\widehat{M})$ is finite.
-Characterize when $M$ is a C-monoid.
-The cases when $M$ is seminormal or strong Mori are of special interest ([74]).
+[…]
 There is a simple characterization of C-monoids within the class of finitely generated monoids and a partial answer for Mori domains. But the general case is open.
@@
   1. (1)
-    Let $M$ be a monoid such that $M_{{\text{\rm red}}}$ is finitely generated.
-Then $M$ is Mori with $(M\negthinspace:\negthinspace\widehat{M})\neq\emptyset$, and $M$ is a C-monoid if and only if $\mathcal{C}_{v}(\widehat{M})$ is finite.
+[…]
   2. (2)
-    Let $D$ be Mori domain with $(D\negthinspace:\negthinspace\widehat{D})\neq\{0\}$ and with $\mathcal{C}_{v}(\widehat{D})$ finite.
-If $\widehat{D}/(D\negthinspace:\negthinspace\widehat{D})$ is finite, then $D$ is a C-domain.
-The converse holds if $D$ is semilocal, Noetherian but not local.
+[…]
 Proof.
-For Part (1), we refer to [66, Proposition 4.8]. The first statement in Part (2) follows from [64] and the converse statement is proved by Reinhart in [127, Corollary 4.5]. ∎
-Every generalized Cohen-Kaplansky domain (this is an atomic domain that has only finitely many atoms, up to associates, that are not prime) is a C-domain ([71, Proposition 4.11]).
+For Part (1), we refer to [68, Proposition 4.8]. The first statement in Part (2) follows from [66] and the converse statement is proved by Reinhart in [128, Corollary 4.5]. ∎
+Every generalized Cohen-Kaplansky domain (this is an atomic domain that has only finitely many atoms, up to associates, that are not prime) is a C-domain ([73, Proposition 4.11]).
 Part 2 of Theorem 12.4 shows that, in particular, orders in algebraic number fields are C-domains.
-A classical result of Furtwängler characterizes the ideals that can be realized as conductor ideals of an order ([87, Theorem 2.12.10]).
+A classical result of Furtwängler characterizes the ideals that can be realized as conductor ideals of an order ([89, Theorem 2.12.10]).
 This was generalized by Reinhart to a more general setting of Noetherian domains ([129]), and his result gives rise to the following problem.
 PROBLEM 19.
-Let $H$ be a Krull monoid with finite $v$-class group.
-Characterize the ideals $\mathfrak{f}\subseteq H$ for which there is a C-monoid $M\subseteq H$ with $\widehat{M}=H$ and with conductor $\mathfrak{f}=(M\negthinspace:\negthinspace\widehat{M})$.
+[…]
 C-domains have been generalized to weakly C-domains which cover a larger class of Mori domains, as given in Theorem 12.4.2.
@@
     If $u\in M$, $b,\,c\in B$ and $\theta(u)=bc$, then there exist $v,\,w\in M$ such that $u=vw$, $\theta(v)\in bB^{\times}$ and $\theta(w)\in cB^{\times}$.
-It is easy to check that $\mathsf{L}_{M}(a)=\mathsf{L}_{B}\big{(}\theta(a)\big{)}$ for all $a\in M$ and $\mathcal{L}(M)=\mathcal{L}(B)$.
-In particular, we have $\rho(M)=\rho(B)$ and $\Delta(M)=\Delta(B)$.
+It is easy to check that $\mathsf{L}_{M}(a)=\mathsf{L}_{B}\big(\theta(a)\big)$ for all $a\in M$ and $\mathcal{L}(M)=\mathcal{L}(B)$. In particular, we have $\rho(M)=\rho(B)$ and $\Delta(M)=\Delta(B)$.
 Let $M$ be a Krull monoid and let $G_{0}\subseteq\mathcal{C}_{v}(M)$ denote the set of classes containing prime divisors.
@@
 Here we restrict our discussion to monoids (so, to commutative cancellative semigroups) and just mention two results.
-A one-dimensional local Mori domain $D$ is transfer Krull if and only if $D$ is half-factorial (whence in the trivial case only; see [73, Theorem 5.5]).
+A one-dimensional local Mori domain $D$ is transfer Krull if and only if $D$ is half-factorial (whence in the trivial case only; see [75, Theorem 5.5]).
 Recent characterizations for orders in Dedekind domains are given in [126].
@@
 For a given class $\mathcal{C}$, for which $\mathcal{V}(\mathcal{C})$ is Krull, study classes $\mathcal{C}^{\prime}$ of modules having a transfer homomorphism $\theta\colon\mathcal{V}(\mathcal{C}^{\prime})\to\mathcal{V}(\mathcal{C})$.
-For a simple (artificial) example, of how this strategy could work, see [75, Example 5.4].
+For a simple (artificial) example, of how this strategy could work, see [77, Example 5.4].
 12.4. What can a family of length sets look like?
-In contrast to Theorem 12.2 there is the following result by Kainrath (see [64, Theorem 7.4.1]).
+In contrast to Theorem 12.2 there is the following result by Kainrath (see [66, Theorem 7.4.1]).
 Theorem 12.5.
@@
 This is a Krull monoid with finite $v$-class group and with a norm function that allows to establish a theory of L-functions ([100]).
-Then almost all (in the sense of Dirichlet density) length sets are intervals ([64, Theorem 9.4.11]).
+Then almost all (in the sense of Dirichlet density) length sets are intervals ([66, Theorem 9.4.11]).
 Other concepts of density include the Zariski density (say for $\mathbb{Q}$-algebras).
@@
 The algebraic and arithmetic structure of ideal semigroups has found wide attention in the literature (for a sample of recent contributions, see [98, 3, 50, 73, 18, 99, 71, 60, 110]).
-In general, these semigroups need not be cancellative (indeed, the semigroup of all nonzero ideals of a domain is cancellative if and only if the domain is almost Dedekind [84, Theorem 23.2]).
+In general, these semigroups need not be cancellative (indeed, the semigroup of all nonzero ideals of a domain is cancellative if and only if the domain is almost Dedekind [86, Theorem 23.2]).
 However, semigroups of ideals are unit-cancellative in a variety of significant cases.